Jefferson Lab HALL A Logbook for December 10 - All EXCEPT Automatic Entries

  • Main INDEX

    Friday

  • 345658 12/31/10 14:35 Jack Carbon-dioxide gas disconnected for eP
  • Thursday

    
    

    Wednesday

    
    

    Monday

  • 345655 12/27/10 21:33 A. Camsonne DVCS Calorimeter LED Low Voltage disconnected
  • 345647 12/27/10 19:57 A. Camsonne DVCS calorimeter HV changed
  • 345646 12/27/10 19:52 A. Camsonne Calorimeter LED scan run 9887
  • 345641 12/27/10 17:23 A. Camsonne LED Low Voltage plugged in for LED run
  • Sunday

  • 345639 12/26/10 01:26 A. Camsonne Trigger threshold set to 10 for run 9885
  • 345636 12/26/10 01:25 A. Camsonne DVCS calorimeter HV increased by 50 V
  • Saturday

  • 345627 12/25/10 15:51 A. Camsonne LED Calorimeter ON at 57 Hz
  • 345626 12/25/10 15:24 A. Camsonne Run 9878 long cosmics run
  • Friday

  • 345624 12/24/10 00:27 Abdurahim Compton lasers are off
  • Thursday

  • 345623 12/23/10 20:00 folts/tech end of run
  • 345622 12/23/10 17:10 A. Camsonne Threshold om sum set to 1536*208+40
  • 345616 12/23/10 09:49 A. Camsonne LED trigger off and HV ON on calorimeter and scintillator paddles for cosmics data taking
  • 345615 12/23/10 09:19 V. Sulkosky Scattering Chamber Vacuum Alarm
  • 345614 12/23/10 09:08 V. Sulkosky LHRS Magnets and VDCs Ramped Down
  • 345613 12/23/10 08:07 Carlos GAME OVER
  • 345611 12/23/10 08:01 bogdanw shift summary
  • 345610 12/23/10 07:55 bogdanw BCM calibration
  • 345605 12/23/10 07:00 Kai target moved to EMPTY
  • 345603 12/23/10 06:38 bogdanw 5 uA on dummy target
  • 345601 12/23/10 06:32 Kai Pan move target to Dummy and set loop2 as park mode
  • 345597 12/23/10 06:00 Carlos Rates and datamon for 9871
  • 345596 12/23/10 05:34 Carlos DC currents run 9871
  • 345595 12/23/10 05:33 Carlos Coda and datamon run 9871
  • 345594 12/23/10 05:32 Carlos Rates run 9871
  • 345593 12/23/10 05:31 Carlos spot++ run 9871
  • 345584 12/23/10 04:18 zana Scaler run 9867
  • 345583 12/23/10 04:17 zana Rocs and RunControl run 9867
  • 345582 12/23/10 04:04 Kai Pan target snapshot
  • 345579 12/23/10 03:19 zana Scalers run 9866
  • 345578 12/23/10 03:18 zana Rocs and RunControl run 9866
  • 345575 12/23/10 02:17 zana Scalers run 9865
  • 345574 12/23/10 02:17 zana Rocs and RunControl run 9865
  • 345573 12/23/10 02:09 bogdanw DVCS ADC
  • 345571 12/23/10 01:20 zana Scalers run 9864
  • 345570 12/23/10 01:20 zana Rocs and RunControl run 9864
  • 345566 12/23/10 00:10 G.M. Urciuol swing shift report
  • Wednesday

  • 345565 12/22/10 23:48 Dmitri target snapshots
  • 345561 12/22/10 23:36 Dmitri daq snapshots, run 9862
  • 345560 12/22/10 23:35 Dmitri scalers and 4A/4B positions, run 9862
  • 345559 12/22/10 22:54 G.M. Urciuol Calorimeter ADC
  • 345556 12/22/10 22:33 Dmitri snapshots, run 9861
  • 345555 12/22/10 22:30 Dmitri scalers run 9861, positions 4A/4B
  • 345554 12/22/10 21:57 Dmitri target snapshots
  • 345553 12/22/10 21:47 A. Camsonne Compton HV Off
  • 345552 12/22/10 21:45 A. Camsonne Compton rates at 1500 V
  • 345551 12/22/10 21:42 A. Camsonne Compton rates high ( as expected )
  • 345550 12/22/10 21:28 A. Camsonne Threshold changed to 70 for DVCSLH2 config
  • 345546 12/22/10 21:22 Dmitri target changed to LH2
  • 345545 12/22/10 21:15 G.M. Urciuol DVCS IOC rebooted
  • 345544 12/22/10 19:56 Dmitri target snapshots
  • 345541 12/22/10 18:30 Dmitri Loop 2 unparked to production
  • 345540 12/22/10 17:56 Dmitri target snapshots
  • 345538 12/22/10 16:10 Dmitri target snapshots
  • 345536 12/22/10 11:16 V. Sulkosky LHRS Set for Kin 1 High Energy
  • 345535 12/22/10 11:12 V. Sulkosky IOC for Magnet Controls rebooted
  • 345534 12/22/10 10:31 A. Camsonne Left HRS angle set to 16.37 degrees
  • 345527 12/22/10 08:04 A. Deur Target alarms due to haH2_P_Balast
  • 345526 12/22/10 07:53 Dmitri target snapshots
  • 345524 12/22/10 06:36 Hisham Beam will be down till after Christmas
  • 345523 12/22/10 05:58 Hisham Machine in down till Christmas
  • 345522 12/22/10 05:56 Dmitri target snapshots
  • 345521 12/22/10 05:11 Dmitri haH2_P_Ballast_R alarm
  • 345520 12/22/10 03:53 Dmitri target snapshots
  • 345514 12/22/10 02:32 Dmitri h2_Ballast_R major alarm
  • 345513 12/22/10 01:58 Dmitri target snapshots
  • 345512 12/22/10 00:47 Hisham Beamm down
  • 345511 12/22/10 00:07 Dmitri target snapshots
  • 345510 12/22/10 00:02 G.M. Urciuol swing shit report
  • Tuesday

  • 345508 12/21/10 23:49 G.M. Urciuo Calorimeter ADC
  • 345506 12/21/10 22:57 itard entries end of run 9855
  • 345505 12/21/10 22:57 itard scaler data run 9855
  • 345504 12/21/10 22:27 G.M. Urciuol DVCS IOC rebooted
  • 345501 12/21/10 21:49 itard scaler data run 9854
  • 345500 12/21/10 21:49 G.M. Urciuol adaql2 before end of run 9854
  • 345499 12/21/10 21:48 G.M. Urciuol hapc5 before end of run 9854
  • 345498 12/21/10 20:51 A. Camsonne LD2 config changed to have a threshold of 80
  • 345497 12/21/10 20:51 G.M. Urciuol CODA condiguration changed
  • 345492 12/21/10 20:46 G.M. Urciuol hapc5 before end of run 9852
  • 345491 12/21/10 20:45 itard entries end of run 9852
  • 345488 12/21/10 19:47 itard entries end of run 9851
  • 345487 12/21/10 19:47 itard scaler data run 9851
  • 345480 12/21/10 18:43 V. Sulkosky Changed CODA Config
  • 345478 12/21/10 17:31 itard entries end of run 9850
  • 345477 12/21/10 17:31 itard scaler data run 9850
  • 345474 12/21/10 16:38 itard All "CALO TRIG" anormalous
  • 345473 12/21/10 16:32 itard entries end of run 9849
  • 345472 12/21/10 16:32 itard scaler data run 9849
  • 345470 12/21/10 15:58 Kai Pan day shift summay
  • 345468 12/21/10 15:35 Kai xscaler for run 9848
  • 345466 12/21/10 15:34 Kai Pan ROC for run 9848
  • 345465 12/21/10 14:42 X. Zheng target status
  • 345463 12/21/10 14:31 Kai Pan scaler for run 9847
  • 345461 12/21/10 14:30 Kai Pan DAQ and ROC for run 9847
  • 345458 12/21/10 13:31 X. Zheng Runcontrol and ROCs run 9846
  • 345457 12/21/10 13:30 X. Zheng scaler run 9846
  • 345456 12/21/10 13:29 X. Zheng Target L2 JT valve closed 2 steps
  • 345454 12/21/10 12:30 X. Zheng runs 9844 and 9845
  • 345452 12/21/10 12:25 X. Zheng Runcontrol and ROCs run 9845
  • 345451 12/21/10 12:24 X. Zheng scaler run 9845
  • 345449 12/21/10 11:52 X. Zheng Runcontrol and ROCs run 9844
  • 345448 12/21/10 11:51 X. Zheng scaler run 9844
  • 345442 12/21/10 11:19 A. Camsonne Compton rates background increased
  • 345441 12/21/10 11:02 X. Zheng target status
  • 345439 12/21/10 10:55 A. Camsonne ROC17 VME ADC got stuck - CODA Ok now
  • 345437 12/21/10 10:40 jpchen Loop 1 temperature readout problem fixed
  • 345435 12/21/10 10:04 X. Zheng Run control and ROC run 9841
  • 345434 12/21/10 10:03 X. Zheng scaler run 9841
  • 345430 12/21/10 08:06 Hisham Owl shift summary
  • 345429 12/21/10 07:53 Dmitri target snapshots
  • 345426 12/21/10 07:27 zana Scaler run 9839
  • 345425 12/21/10 07:26 zana Rocs and RunControl run 9839
  • 345422 12/21/10 06:22 zana Scaler run 9838
  • 345421 12/21/10 06:22 zana Rocs and RunControl run 9838
  • 345420 12/21/10 06:02 Dmitri target snapshots
  • 345418 12/21/10 05:28 zana Rocs and RunControl run 9837
  • 345416 12/21/10 04:58 hisham DVCS DC levels
  • 345415 12/21/10 04:29 Dmitri target snapshots
  • 345406 12/21/10 02:37 Carlos Switched to adaql2
  • 345405 12/21/10 02:27 Hisham Change configuration to DVCSLH2
  • 345402 12/21/10 02:19 zana Scaler Run 9834
  • 345401 12/21/10 02:18 zana Rocs and RunControl run 9834
  • 345400 12/21/10 02:13 Dmitri target snapshots
  • 345399 12/21/10 01:25 Hisham IOC reboot
  • 345394 12/21/10 01:00 zana Scaler run 9832
  • 345393 12/21/10 00:59 zana Rocs and RunControl run 9832
  • 345392 12/21/10 00:47 jpchen Loop 1 (LD2) temperature sensor problem, PID off, TO notice
  • 345391 12/21/10 00:23 Kai Pan BTW, no alarm sound for target
  • 345390 12/21/10 00:22 Kai Pan target loop1 sensor problem
  • 345389 12/21/10 00:06 bogdanw shift summary
  • Monday

  • 345385 12/20/10 23:59 zana Scaler run 9831
  • 345384 12/20/10 23:58 zana Rocs and RunControl run 9831
  • 345383 12/20/10 23:49 jroche Compton results
  • 345382 12/20/10 23:47 kai pan loop1 diagram
  • 345381 12/20/10 23:21 Kai Pan target loop.1 temp fluctuate
  • 345378 12/20/10 22:58 rakithab CODA status during the run 9830
  • 345377 12/20/10 22:58 rakithab Hall A scaler data during the run 9830
  • 345374 12/20/10 22:01 rakithab CODA status during the run 9829
  • 345373 12/20/10 22:01 rakithab Hall A scaler data during the run 9829
  • 345372 12/20/10 21:54 rakithab Calorimeter ADCl levels
  • 345371 12/20/10 21:15 Kai Pan H2 target ready
  • 345369 12/20/10 21:01 Kai Pan move target to H2
  • 345368 12/20/10 21:00 rakithab CODA configuration Ben1test reloaded for the LH2 target data
  • 345366 12/20/10 20:58 rakithab CODA status during the run 9828
  • 345365 12/20/10 20:57 rakithab Hall A scaler data during the run 9828
  • 345362 12/20/10 20:02 rakithab CODA status during the run 9827
  • 345361 12/20/10 20:01 rakithab Hall A scaler data status during the run 9827
  • 345353 12/20/10 18:57 rakithab CODA status during the run 9826
  • 345352 12/20/10 18:56 rakithab Hall A scaler data during the run 9826
  • 345350 12/20/10 18:02 Kai Pan target snapshot
  • 345348 12/20/10 17:59 rakithab CODA status during the run 9825
  • 345347 12/20/10 17:59 rakithab Hall A scaler data during the run 9825
  • 345346 12/20/10 17:52 A. Camsonne Turned cleanup script temporary off on adaql2
  • 345345 12/20/10 17:05 bogdanw DVCS calo currents
  • 345344 12/20/10 17:04 A. Camsonne Run 9825 running on adaql1
  • 345342 12/20/10 16:49 rakithab CODA status during the run 9824
  • 345341 12/20/10 16:49 rakithab Hall A scaler data during the run 9824
  • 345340 12/20/10 16:13 rakithab CODA (ER) warnings durrng the run 9824
  • 345339 12/20/10 16:01 G.M. Urciuol day shift report
  • 345336 12/20/10 15:54 efuchey Compton coda crash: Complementary information
  • 345335 12/20/10 15:54 A. Camsonne adaql2 disks full
  • 345334 12/20/10 15:51 efuchey Compton IOC reboot
  • 345333 12/20/10 15:45 efuchey Compton coda crash
  • 345332 12/20/10 15:43 Carlos Run 9822 : ROC17 crashed
  • 345330 12/20/10 15:04 efuchey Target charts
  • 345329 12/20/10 15:03 efuchey target screenshots
  • 345328 12/20/10 15:03 G.M. Urciuol DVCS IOC rebooted
  • 345325 12/20/10 14:53 pzhu scaler9821
  • 345324 12/20/10 14:53 pzhu coda9821
  • 345323 12/20/10 14:43 efuchey quarter wave plate ready to go
  • 345322 12/20/10 14:26 G.M. Urciuol Calorimeter ADC
  • 345321 12/20/10 14:22 efuchey QWP stuck.
  • 345320 12/20/10 14:15 efuchey Misplacement of quarter wave plate
  • 345317 12/20/10 13:59 pzhu scaler screen for run9820
  • 345316 12/20/10 13:59 pzhu coda screen for run9820
  • 345315 12/20/10 13:05 efuchey Target charts
  • 345314 12/20/10 13:04 efuchey Target screenshots
  • 345311 12/20/10 12:56 yez Scaler screen for run9819
  • 345310 12/20/10 12:56 yez CODA screens for run#9819
  • 345307 12/20/10 12:01 yez Scaler screen for run9818
  • 345306 12/20/10 12:00 yez CODA screen for run9818
  • 345300 12/20/10 11:29 efuchey reopened J/T valve for loop 1
  • 345299 12/20/10 11:26 Carlos New trigger plots added to calo GUI
  • 345298 12/20/10 11:18 efuchey Target charts
  • 345297 12/20/10 11:18 efuchey Target screenshots
  • 345294 12/20/10 10:48 yez Scaler screens for run9817
  • 345293 12/20/10 10:47 yez CODA screens for run9817
  • 345292 12/20/10 10:40 efuchey J/T valve for loop 2 closed.
  • 345291 12/20/10 10:09 efuchey J/T valve for loop 2
  • 345290 12/20/10 10:02 efuchey J/T valve closed for loop 1
  • 345288 12/20/10 09:57 yez coda screens
  • 345286 12/20/10 09:35 efuchey Target charts
  • 345285 12/20/10 09:32 efuchey Target screenshots
  • 345282 12/20/10 08:47 yez Scaler screens
  • 345281 12/20/10 08:47 yez CODA screens
  • 345278 12/20/10 08:25 yez Scaler screen snapshots
  • 345277 12/20/10 08:24 yez CODA screen snapshots
  • 345276 12/20/10 08:06 Hisham Owl shift summary
  • 345274 12/20/10 07:46 Dmitri target snapshots
  • 345271 12/20/10 07:28 itard entries end of run 9813
  • 345270 12/20/10 07:28 itard scaler data run 9813
  • 345267 12/20/10 06:31 itard entries end of run 9812
  • 345266 12/20/10 06:31 itard scaler data run 9812
  • 345265 12/20/10 06:13 Dmitri target snapshots
  • 345263 12/20/10 05:32 itard entries end of run 9811
  • 345262 12/20/10 05:32 itard scaler data run 9811
  • 345261 12/20/10 04:43 Hisham DVCS DC levels
  • 345259 12/20/10 04:30 itard entries end of run 9810
  • 345258 12/20/10 04:30 itard scaler data run 9810
  • 345257 12/20/10 04:15 Dmitri target snapshots
  • 345249 12/20/10 03:31 itard entries end of run 9809
  • 345248 12/20/10 03:31 itard scaler data run 9809
  • 345245 12/20/10 02:30 itard entries end of run 9808
  • 345244 12/20/10 02:30 itard scaler data run 9808
  • 345243 12/20/10 02:19 Dmitri target snapshots
  • 345240 12/20/10 01:31 itard entry end of run 9807
  • 345239 12/20/10 01:31 itard scaler data run 9807
  • 345236 12/20/10 00:33 itard entries end of run 9806
  • 345235 12/20/10 00:33 itard scaler data run 9806
  • 345234 12/20/10 00:12 Dmitri target snapshots
  • 345232 12/20/10 00:01 LeRose Shift Summary (swing)
  • Sunday

  • 345229 12/19/10 23:15 garibaldi scalers run 9805
  • 345228 12/19/10 23:14 garibaldi rocs run 9805
  • 345227 12/19/10 22:37 D.Wang Target temperature drop
  • 345226 12/19/10 22:28 Abdurahim Compton Status
  • 345223 12/19/10 22:10 garibaldi scalers run 9804
  • 345222 12/19/10 22:09 garibaldi rocs run 9804
  • 345219 12/19/10 21:18 garibaldi scalers run 9803
  • 345218 12/19/10 21:17 garibaldi rocs run 9803
  • 345217 12/19/10 21:13 D.Wang Target Status
  • 345216 12/19/10 20:37 R. Michaels more on disk management.
  • 345215 12/19/10 20:34 LeRose Calorimeter ADC
  • 345213 12/19/10 20:18 garibaldi scalers run 9802
  • 345212 12/19/10 20:17 garibaldi rocs run 9802
  • 345206 12/19/10 19:33 LeRose DVCS IOC reboot
  • 345200 12/19/10 18:17 garibaldi scalers run 9800
  • 345199 12/19/10 18:17 garibaldi rocs run 9800
  • 345198 12/19/10 18:07 D.Wang Target Status
  • 345197 12/19/10 17:21 Abdurahim Compton cavity has been restored
  • 345195 12/19/10 17:19 garibaldi scalers run 9799
  • 345194 12/19/10 17:18 garibaldi rocs run 9799
  • 345193 12/19/10 17:15 D.Wang target alarm haH2_P_Ballast_R
  • 345192 12/19/10 17:10 R. Michaels increased frequency of copying
  • 345191 12/19/10 16:46 R. Michaels increased frequency of cleanup
  • 345190 12/19/10 16:39 R. Michaels checks of tape writing
  • 345189 12/19/10 16:23 V. Sulkosky CODA Restarted 3 times
  • 345188 12/19/10 16:22 D.Wang Target Status
  • 345187 12/19/10 16:22 efuchey Problems with cavity
  • 345183 12/19/10 16:09 Carlos Disks full: writing to adaql1
  • 345181 12/19/10 16:07 G.M. Urciuol day shift report
  • 345179 12/19/10 15:13 efuchey Target change
  • 345175 12/19/10 14:56 G.M. Urciuol adaql2 before end of run 9794
  • 345174 12/19/10 14:55 G.M. Urciuol hapc5 before end of run 9794
  • 345173 12/19/10 13:59 G.M. Urciuol Calorimeter ADC
  • 345170 12/19/10 13:42 G.M. Urciuol hapc5 before end of run 9793
  • 345169 12/19/10 13:40 G.M. Urciuol adaql2 before end of run 9793
  • 345168 12/19/10 13:11 efuchey Target charts
  • 345167 12/19/10 13:09 efuchey Target screenshots
  • 345164 12/19/10 12:41 efuchey Compton run 24693
  • 345163 12/19/10 12:36 G.M. Urciuol adaql2 before end of run 9792
  • 345162 12/19/10 12:35 G.M. Urciuol hapc5 before end of run 9792
  • 345161 12/19/10 11:47 efuchey JT Valve close
  • 345153 12/19/10 11:35 G.M. Urciuol adaql2 before end of run 9791
  • 345152 12/19/10 11:33 G.M. Urciuol hapc5 before end of run 9791
  • 345151 12/19/10 11:20 efuchey Compton run 24691
  • 345150 12/19/10 11:01 efuchey Target charts
  • 345149 12/19/10 10:59 efuchey Back to compton production
  • 345148 12/19/10 10:56 efuchey Target screenshots
  • 345147 12/19/10 10:50 efuchey QWP scan for 12/19/2010
  • 345144 12/19/10 10:33 G.M. Urciuol adaql2 before 9790 end of run
  • 345143 12/19/10 10:32 efuchey scaler rates run 9790
  • 345142 12/19/10 10:02 G.M. Urciuol DVCS IOC rebooted
  • 345140 12/19/10 09:33 efuchey Lumis behavior (again)
  • 345138 12/19/10 09:26 efuchey Scaler rates for run 9789
  • 345137 12/19/10 09:26 G.M. Urciuol adaql2 Before End Of Run 9789
  • 345136 12/19/10 09:15 efuchey Compton stopped
  • 345135 12/19/10 09:06 efuchey Target Charts
  • 345132 12/19/10 09:00 efuchey Target screenshots
  • 345131 12/19/10 08:58 G.M. Urciuol adaql2 before end of run 9788
  • 345130 12/19/10 08:57 G.M. Urciuol hapc5 before End Of Run 9788
  • 345129 12/19/10 08:55 efuchey ROC 4 and ROC 14
  • 345128 12/19/10 08:46 efuchey lumis behavior
  • 345127 12/19/10 08:07 Hisham Owl shift summary
  • 345123 12/19/10 07:58 buddhini Scaler rates before end of run 9787
  • 345122 12/19/10 07:58 buddhini CODA windows before end of run 9787
  • 345121 12/19/10 07:14 Dmitri target snapshots
  • 345118 12/19/10 07:01 buddhini CODA windows before the end of run 9786
  • 345117 12/19/10 07:00 buddhini Scaler rates before the end of run 9786
  • 345114 12/19/10 06:05 buddhini Scaler rates before the end of run 9785
  • 345113 12/19/10 06:04 buddhini CODA windows before end of run 9785
  • 345110 12/19/10 05:04 Dmitri target snapshots
  • 345109 12/19/10 04:59 buddhini Scaler rates before the end of run 9784
  • 345108 12/19/10 04:59 buddhini CODA screens before end of run 9784
  • 345107 12/19/10 04:55 Hisham DVCS DC levels
  • 345106 12/19/10 04:45 Hisham Cavity didn't lock for 20 minutes
  • 345103 12/19/10 04:02 buddhini Scaler rates before end of run 9783
  • 345102 12/19/10 04:01 buddhini CODA windows before end of run 9783
  • 345096 12/19/10 03:42 Hisham Compton End run screen
  • 345093 12/19/10 03:04 buddhini CODA windows before end of run 9782
  • 345092 12/19/10 03:04 buddhini Scaler rates before end of run 9782
  • 345091 12/19/10 03:04 Dmitri target snapshots
  • 345090 12/19/10 02:07 Hisham IOC reboot during run 9781
  • 345087 12/19/10 02:04 buddhini CODA windows at the end of run 9781
  • 345084 12/19/10 01:37 buddhini Scaler rates before end of run 9780
  • 345083 12/19/10 01:37 buddhini CODA windows before end of run 9780
  • 345082 12/19/10 01:16 Dmitri target snapshots
  • 345080 12/19/10 00:40 buddhini Scaler rates before end of run 9779
  • 345078 12/19/10 00:39 buddhini CODA screens beofore end of run 9779
  • Saturday

  • 345076 12/18/10 23:55 eschulte Swing Shift Log
  • 345075 12/18/10 23:55 eschulte Swing Shift Runs
  • 345073 12/18/10 23:38 rakithab Hall A scaler status during the run 9778
  • 345072 12/18/10 23:38 rakithab CODA status during the run 9778
  • 345070 12/18/10 23:01 siqbal target status
  • 345069 12/18/10 22:59 siqbal JP changed fan settings (troubleshooting)
  • 345066 12/18/10 22:28 rakithab CODA status during the run 9777
  • 345065 12/18/10 22:26 rakithab Hall A scaler data during the run 9777
  • 345064 12/18/10 21:40 siqbal target status
  • 345061 12/18/10 21:34 rakithab CODA status during the run 9776
  • 345060 12/18/10 21:32 rakithab Hall A scaler data during the run 9776
  • 345059 12/18/10 21:22 eschulte Compton Strip Chart: Last hour
  • 345056 12/18/10 20:30 rakithab CODA status during the run 9775
  • 345055 12/18/10 20:29 rakithab Hall A scaler data during the run 9775
  • 345054 12/18/10 20:26 eschulte Swing Shift Checklist: Calo Currents
  • 345048 12/18/10 19:41 siqbal target status
  • 345045 12/18/10 19:27 rakithab CODA status during the run 9774
  • 345044 12/18/10 19:26 rakithab Hall A scaler data during the run 9774
  • 345043 12/18/10 18:49 eschulte Compton: Reboot ROC6 no effect on ER
  • 345040 12/18/10 18:27 rakithab CODA status during the run 9773
  • 345039 12/18/10 18:26 rakithab Hall A Scaler data during the run 9773
  • 345038 12/18/10 17:47 eschulte Compton: DAQ Troubles
  • 345037 12/18/10 17:46 V. Sulkosky Corrected ABUs for Day Shift 12/18/2010
  • 345036 12/18/10 17:33 siqbal target status
  • 345034 12/18/10 17:23 rakithab Hall A scaler data during the run 9772
  • 345033 12/18/10 17:22 rakithab CODA status during the run 9772.
  • 345032 12/18/10 16:34 R. Michaels MSS back up, copying automated again
  • 345029 12/18/10 16:24 rakithab CODA Status during the run 9771.
  • 345028 12/18/10 16:24 rakithab Hall A scaler data status during the run 9771
  • 345027 12/18/10 16:10 G.M. Urciuol day shift report
  • 345025 12/18/10 15:45 efuchey Loop 1 fan frequency stuck.
  • 345023 12/18/10 15:27 efuchey Target Charts
  • 345022 12/18/10 15:23 efuchey JT valves
  • 345021 12/18/10 15:16 efuchey Target move to LH2
  • 345019 12/18/10 14:58 pzhu coda9770
  • 345018 12/18/10 14:58 pzhu scaler9770
  • 345017 12/18/10 14:28 G.M. Urciuol Calorimeter ADC
  • 345016 12/18/10 14:23 Carlos Rates fluctuations
  • 345011 12/18/10 13:59 pzhu scaler9768
  • 345010 12/18/10 13:59 pzhu coda9768
  • 345009 12/18/10 13:17 G.M. Urciuol Fluctuations in deadtime
  • 345008 12/18/10 13:02 efuchey Target charts
  • 345006 12/18/10 13:00 efuchey Target screenshots
  • 345004 12/18/10 12:58 pzhu scaler9767
  • 345003 12/18/10 12:58 pzhu coda9767
  • 345002 12/18/10 12:19 R. Michaels modified end_clean
  • 344999 12/18/10 11:48 pzhu scaler9766
  • 344998 12/18/10 11:48 pzhu coda9766
  • 344992 12/18/10 11:26 A. Camsonne Compton hac_bcm_average timeout problem
  • 344991 12/18/10 11:20 R. Michaels data movements without MSS
  • 344990 12/18/10 11:10 efuchey Target Charts
  • 344989 12/18/10 11:08 efuchey Target screenshots
  • 344986 12/18/10 10:48 pzhu scaler9765
  • 344985 12/18/10 10:46 pzhu coda9765
  • 344984 12/18/10 10:36 efuchey Can't get rid of flatlining rates
  • 344983 12/18/10 10:36 A. Camsonne New configuration for data taking
  • 344982 12/18/10 10:07 efuchey CODA crash
  • 344981 12/18/10 09:58 efuchey Compton CODA froze again.
  • 344978 12/18/10 09:31 A. Camsonne Power cycled all crates DAQ fine now
  • 344975 12/18/10 09:22 A. Camsonne Network still slow after power cycling the two network switches
  • 344974 12/18/10 09:21 efuchey Target charts
  • 344973 12/18/10 09:09 efuchey Poltarac reboooted
  • 344972 12/18/10 09:07 efuchey Target screenshots
  • 344971 12/18/10 08:09 Hisham Owl shift summary
  • 344969 12/18/10 07:40 buddhini CODA windows before end of run 9762. CODA froze.
  • 344966 12/18/10 07:05 Dmitri target snapshots
  • 344965 12/18/10 07:05 buddhini Scaler rates before end of run 9761
  • 344964 12/18/10 07:04 buddhini CODA windos before end of run 9761
  • 344963 12/18/10 06:17 Hisham Compton HV OFF
  • 344962 12/18/10 06:16 Hisham Compton problem persist
  • 344961 12/18/10 06:08 buddhini Scaler rates before the end of run 9760
  • 344959 12/18/10 06:07 buddhini CODA windows at the end of run 9760
  • 344957 12/18/10 05:58 Hisham DVCS DC levels
  • 344956 12/18/10 05:45 Hisham DVCS HV trip during run 9759
  • 344955 12/18/10 05:19 Dmitri target snapshots
  • 344954 12/18/10 05:10 Hisham DVCS high voltage trip
  • 344951 12/18/10 04:55 Hisham Coda reboot (3rd time)
  • 344948 12/18/10 04:51 buddhini CODA windows at the end of run 9758
  • 344946 12/18/10 04:24 Hisham Compton problems
  • 344944 12/18/10 04:12 buddhini Scaler rates before end of run 9757
  • 344943 12/18/10 04:12 Hisham Compton Coda reboot (2nd time)
  • 344942 12/18/10 04:12 buddhini CODA windows before end of run 9757
  • 344940 12/18/10 04:07 Hisham Compton rates flat
  • 344939 12/18/10 03:58 Hisham Coda reboot for Compton
  • 344933 12/18/10 03:20 R. Michaels plan for MSS outage
  • 344930 12/18/10 03:07 buddhini Scaler rates before the end of run 9756
  • 344929 12/18/10 03:06 buddhini CODA screens before end of run 9756
  • 344928 12/18/10 03:02 Dmitri target snapshots
  • 344926 12/18/10 02:08 buddhini Scaler rates befoe the end of run 9755
  • 344925 12/18/10 02:07 adaq CODA windows before end of run 9756
  • 344923 12/18/10 01:15 Dmitri target snapshots
  • 344922 12/18/10 01:11 buddhini Scaler rates before end of run 9754
  • 344920 12/18/10 01:10 buddhini CODA screens before end of run 9754
  • 344919 12/18/10 00:07 buddhini scaler rates at end of run 9753
  • 344917 12/18/10 00:06 adaq CODA windows at end of run 9753
  • 344914 12/18/10 00:02 eschulte Swing Shift Log
  • Friday

  • 344913 12/17/10 23:44 eschulte Swing Shift Runs
  • 344910 12/17/10 23:08 Aniol rocs 9752
  • 344909 12/17/10 23:07 Aniol coda r9752
  • 344908 12/17/10 23:07 Aniol scalers, r9752
  • 344907 12/17/10 22:54 iqbal target status
  • 344904 12/17/10 22:12 Aniol coda r9751
  • 344903 12/17/10 22:12 Aniol rocs r9751
  • 344902 12/17/10 22:11 Aniol scalers r9751
  • 344901 12/17/10 21:45 Carlos New plot added to calo online GUI
  • 344900 12/17/10 21:41 Carlos How to access golden file from a macro
  • 344895 12/17/10 20:40 eschulte MCC: Manget Trip
  • 344894 12/17/10 20:33 eschulte Swing Shift Checklist calorimeter currents
  • 344891 12/17/10 20:04 Aniol rocs r9748
  • 344890 12/17/10 20:04 Aniol coda r9748
  • 344889 12/17/10 20:03 Aniol scalers r9748
  • 344888 12/17/10 19:55 siqbal target status
  • 344880 12/17/10 19:03 Aniol coda r9747
  • 344879 12/17/10 19:02 Aniol rocs r9747
  • 344878 12/17/10 19:01 Aniol scalers r9747
  • 344877 12/17/10 18:07 A. Camsonne Run 9747 back on adaql2
  • 344874 12/17/10 18:04 Aniol scalers r9746
  • 344873 12/17/10 18:03 Aniol rocs r4796
  • 344872 12/17/10 18:02 Aniol coda r9746
  • 344871 12/17/10 17:41 siqbal target status
  • 344869 12/17/10 17:02 A. Camsonne Set singles prescales factors back to 0 for run 9746
  • 344867 12/17/10 16:58 R. Michaels data file deletion script made more aggressive now
  • 344866 12/17/10 16:38 A. Camsonne Run 9745 on adaql1 - with some singles triggers
  • 344864 12/17/10 16:34 A. Camsonne Test singles triggers
  • 344861 12/17/10 16:29 aniol coda
  • 344860 12/17/10 16:28 aniol scalers r9743
  • 344859 12/17/10 16:24 Aniol rocs+coda r9743
  • 344858 12/17/10 16:07 Holmstrom Day SHift Summary
  • 344856 12/17/10 15:50 Fansler Moeller
  • 344855 12/17/10 15:36 Dmitri target charts snapshot
  • 344854 12/17/10 15:35 Dmitri target gui snapshot
  • 344851 12/17/10 15:20 garibaldi scalers run 9742
  • 344850 12/17/10 15:20 garibaldi rocs 9742
  • 344849 12/17/10 14:59 A. Camsonne ARSTrigger.h and dvcstlab6 bootscript modified
  • 344846 12/17/10 14:23 garibaldi scalers run 9741
  • 344845 12/17/10 14:22 garibaldi Rocs run 9741
  • 344844 12/17/10 14:21 Carlos Writing to adaql1
  • 344843 12/17/10 14:16 Dmitri target charts snapshot
  • 344842 12/17/10 14:15 Dmitri target snapshot main gui
  • 344841 12/17/10 13:59 jpchen Loop2 (LH2) also cooled.
  • 344840 12/17/10 13:24 Holmstrom DVCS IOC reboot
  • 344837 12/17/10 13:07 Dmitri charts screen
  • 344836 12/17/10 13:05 Carlos Run 9739: half optics/half LD2
  • 344833 12/17/10 12:54 jpchen Loop 1 (LD2) cooled down to operational condition
  • 344831 12/17/10 12:21 jpchen Cryo 15 K supply temperature
  • 344830 12/17/10 12:14 Holmstrom Target Temp Flow
  • 344828 12/17/10 11:47 Dmitri hal2_Cemox_1_R and CFI6731C dropped
  • 344816 12/17/10 10:32 garibaldi scalers run 9732
  • 344815 12/17/10 10:31 garibaldi rocs run 9732
  • 344814 12/17/10 10:13 Dmitri main gui screen
  • 344813 12/17/10 10:11 Dmitri parked Loop2 heater dropped to zero
  • 344812 12/17/10 09:44 A. Camsonne Pion prescale set to 200
  • 344811 12/17/10 09:37 Carlos T6 and T5 rates when threshold change
  • 344810 12/17/10 09:35 Dmitri hal2_Cernox_2_R back to normal
  • 344807 12/17/10 09:33 A. Camsonne Trigger threshold set to 80 for deuterium running
  • 344805 12/17/10 09:28 Dmitri charts after target change
  • 344804 12/17/10 09:26 Dmitri target change
  • 344801 12/17/10 09:01 garibaldi scalers run 9730
  • 344800 12/17/10 09:01 Dmitri snapshot of main picture gui
  • 344799 12/17/10 09:01 garibaldi rocs run 9730
  • 344798 12/17/10 08:59 Dmitri target: drop of hal2 value
  • 344795 12/17/10 08:05 garibaldi scalers run 9729
  • 344794 12/17/10 08:04 garibaldi rocs run 9729
  • 344793 12/17/10 08:02 Kai Pan scaler for Run 9729
  • 344792 12/17/10 08:02 R. Michaels simple temp. soln. to missing end-run
  • 344791 12/17/10 08:02 Kai Pan DAQ and ROC for Run 9729
  • 344789 12/17/10 08:00 G. M. Urciuo Owl shift report
  • 344788 12/17/10 07:29 R. Michaels missing halog end-run problem
  • 344785 12/17/10 07:02 Kai Pan scaler for run 9728
  • 344784 12/17/10 07:02 Kai Pan ROC and DAQ for Run 9728
  • 344782 12/17/10 05:12 G.M. Ucciuol Calorimiter ADC
  • 344781 12/17/10 05:06 Kai Pan Run 9726 coda crashed
  • 344779 12/17/10 05:04 Kai Pan scaler Run 9726
  • 344778 12/17/10 05:03 Kai Pan ROC and DAQ for Run 9726 (coda crashed)
  • 344770 12/17/10 03:03 R. Michaels back to adaql2 disks
  • 344768 12/17/10 02:56 Kai Pan scaler Run 9724
  • 344767 12/17/10 02:55 Kai Pan 9724 ROC and DAQ
  • 344766 12/17/10 02:44 Kai Pan rootfile diskfull
  • 344764 12/17/10 01:55 R. Michaels disk space filled up / CODA crash
  • 344761 12/17/10 01:12 Kai Pan scaler for run 9722
  • 344760 12/17/10 01:12 Kai Pan ROC and DAQ for Run 9722
  • 344756 12/17/10 00:05 bogdanw shift summary
  • Thursday

  • 344753 12/16/10 23:14 O.Glamazdin ROCs and DAQ for run 9720
  • 344752 12/16/10 23:13 O.Glamazdin Scalers for run 9720
  • 344749 12/16/10 22:14 O.Glamazdin ROCs and DAQ for run 9719
  • 344748 12/16/10 22:13 O.Glamazdin Scalers for run 9719
  • 344747 12/16/10 21:28 Aniol compton DAQ is slow
  • 344745 12/16/10 21:19 O.Glamazdin CODA crush
  • 344744 12/16/10 21:12 Aniol target screens
  • 344743 12/16/10 21:03 O.Glamazdin ROCs and DAQ for run 9718
  • 344742 12/16/10 21:02 O.Glamazdin Scalers for run 9718
  • 344741 12/16/10 20:54 bogdanw DVCS currents
  • 344738 12/16/10 19:58 O.Glamazdin ROCs and DAQ for run 9717
  • 344737 12/16/10 19:58 O.Glamazdin Scalers for run 9717
  • 344727 12/16/10 18:43 O.Glamazdin Scalers for run 9715
  • 344725 12/16/10 18:40 O.Glamazdin ROCs and DAQ screens for Bogdan's run (9715)
  • 344724 12/16/10 18:38 Aniol target moved to tantalum
  • 344722 12/16/10 18:19 O.Glamazdin ROCs and DAQ for run 9714
  • 344720 12/16/10 18:15 O.Glamazdin DVCS IOC trip
  • 344716 12/16/10 17:23 O.Glamazdin ROCs and DAQ for run 9712
  • 344715 12/16/10 17:22 O.Glamazdin Scalers for run 9712
  • 344714 12/16/10 16:55 Aniol target - set Loop1 to 24Hz
  • 344712 12/16/10 16:36 bogdanw Hall A tools
  • 344710 12/16/10 16:12 O.Glamazdin ROCs and DAQ for run 9711
  • 344709 12/16/10 16:11 O.Glamazdin Scalers for run 9711
  • 344707 12/16/10 15:59 efuchey 12-16-10 day shift summary
  • 344702 12/16/10 14:51 efuchey Calo ADC currents
  • 344701 12/16/10 14:50 efuchey Hall A tools
  • 344700 12/16/10 14:49 Hisham Calo ADC
  • 344698 12/16/10 14:23 hisham Scalers for run 9708
  • 344697 12/16/10 14:22 Hisham Coda, ROCs for run 9708
  • 344696 12/16/10 13:37 hisham target status
  • 344693 12/16/10 13:16 hisham Scalers run 9707
  • 344692 12/16/10 13:15 Hisham ROCs & Coda for run 9707
  • 344690 12/16/10 12:07 X. Zheng Target status
  • 344683 12/16/10 10:20 A. Camsonne Calorimeter trigger firmware downloaded version december 16th for run 9706
  • 344680 12/16/10 10:17 A. Camsonne adaqfs UPS seems dead - bypassed it for now
  • 344679 12/16/10 09:31 A. Camsonne Threshold and pedestal updated - will be taken into account at next download
  • 344678 12/16/10 09:22 X. Zheng Target loop 1 put to sleep
  • 344675 12/16/10 09:14 X. Zheng Target moved to 15cm LH2
  • 344672 12/16/10 09:09 Carlos Rates run 9703
  • 344671 12/16/10 09:08 Carlos ROCs run 9703
  • 344670 12/16/10 09:06 X. Zheng Target loop 2 ready for beam
  • 344669 12/16/10 08:46 A. Camsonne Run taken with adaql1
  • 344666 12/16/10 08:06 A. Camsonne deadtime raised to 30 % turning adaql1 writing on to check it
  • 344664 12/16/10 08:04 G.M. Urciuol owl shift report
  • 344663 12/16/10 07:59 G.M. Urciuol Compton CODA rebooted
  • 344662 12/16/10 07:48 G.M. Urciuol Calorimeter ADC
  • 344659 12/16/10 07:37 itard entries end of run 9701
  • 344658 12/16/10 07:37 itard data scaler run 9701
  • 344657 12/16/10 07:31 G.M. Urciuol Wrong or missing EOR files
  • 344654 12/16/10 06:32 itard entries end of run 9700
  • 344653 12/16/10 06:32 itard scaler data run 9700
  • 344652 12/16/10 06:04 siqbal target status
  • 344650 12/16/10 05:32 itard scaler data run 9699
  • 344649 12/16/10 05:32 itard entries end of run 9699
  • 344646 12/16/10 04:32 itard entries end of run 9698
  • 344645 12/16/10 04:32 itard data scaler run 9898
  • 344644 12/16/10 04:01 siqbal target status
  • 344643 12/16/10 03:48 A. Camsonne modified runend.tcl to allow running on adaql1
  • 344637 12/16/10 03:38 A. Camsonne L1A fluctuations gone even writing on adaql1
  • 344636 12/16/10 03:36 A. Camsonne switching back to adaql2
  • 344633 12/16/10 03:27 itard data scaler run 9697
  • 344632 12/16/10 03:26 itard entries end of run 9693
  • 344631 12/16/10 03:12 A. Camsonne data looks ok using adaql1
  • 344630 12/16/10 02:56 A. Camsonne Random trigger not working
  • 344629 12/16/10 02:42 A. Camsonne Network ok now
  • 344628 12/16/10 02:38 A. Camsonne Test config BenL1test
  • 344624 12/16/10 02:14 R. Michaels re: deadtime / adaql1
  • 344623 12/16/10 02:10 R. Michaels ok, I see why alarm disabled
  • 344621 12/16/10 02:09 siqbal target status
  • 344620 12/16/10 02:01 A. Camsonne Large dead time fluctuation most likely due to adaql1 writing
  • 344619 12/16/10 02:00 R. Michaels DAQ recovery. Disks filled up. Using adaql1 disks
  • 344617 12/16/10 01:22 Carlos Calorimeter golden run updated to 9694
  • 344616 12/16/10 01:18 Carlos DAQ alarm disabled
  • 344614 12/16/10 00:57 itard entries end of run 9693
  • 344613 12/16/10 00:57 itard scaler data run 9693
  • 344612 12/16/10 00:17 A. Camsonne trigger gate width set to 17 - will be taken into account at next download run 9694
  • 344611 12/16/10 00:10 Aniol shift summary 2nd try
  • 344610 12/16/10 00:02 Aniol shift summary
  • Wednesday

  • 344605 12/15/10 23:52 O.Glamazdin DVCS IOC crush.
  • 344602 12/15/10 23:17 A. Camsonne ROC network very slow - power cycling network switch
  • 344601 12/15/10 23:04 O.Glamazdin ROCs and DAQ screens for run 9690
  • 344600 12/15/10 23:04 O.Glamazdin ROCs and DAQ for run 9690
  • 344599 12/15/10 23:03 O.Glamazdin Scalers for run 9690
  • 344598 12/15/10 22:43 Carlos T5 rates
  • 344597 12/15/10 22:29 A. Camsonne trigger gate width set to 15 - will be taken into account at next download
  • 344596 12/15/10 22:29 Carlos T6 rates at 1.995 uA and Gate width 100 ns
  • 344594 12/15/10 21:56 A. Camsonne Trigger firmware Dec 15th loaded
  • 344593 12/15/10 21:55 A. Camsonne Added 22 ns of delay on S2m at the trigger module
  • 344591 12/15/10 21:16 O.Glamazdin ROCs and DAQ screens for run 9689
  • 344590 12/15/10 21:15 O.Glamazdin ROCs and DAQ screens for run 9689
  • 344589 12/15/10 21:13 O.Glamazdin Scalers for run 9689
  • 344588 12/15/10 20:44 Aniol calo ADC readbacks
  • 344587 12/15/10 20:30 O.Glamazdin Run 9688 - wrong CODA configuration
  • 344583 12/15/10 20:21 O.Glamazdin DAQ recovery.
  • 344582 12/15/10 20:11 O.Glamazdin DATAMON problem
  • 344581 12/15/10 20:10 O.Glamazdin CODA crush
  • 344574 12/15/10 19:34 O.Glamazdin Run 9684 was ended due to DVCS IOC crush.
  • 344570 12/15/10 19:00 Aniol target screens
  • 344569 12/15/10 18:59 O.Glamazdin ROCs and DAQ screens for run 9683
  • 344568 12/15/10 18:59 O.Glamazdin ROCs and DAQ screens for run 9683
  • 344567 12/15/10 18:58 O.Glamazdin Scalers for run 9683
  • 344564 12/15/10 17:59 O.Glamazdin ROCs and DAQ screens for run 9682
  • 344563 12/15/10 17:58 O.Glamazdin ROCs and DAQ screens for run 9682
  • 344562 12/15/10 17:58 O.Glamazdin Scalers for run 9682
  • 344559 12/15/10 16:59 O.Glamazdin ROCs and DAQ screens for run 9681
  • 344558 12/15/10 16:58 O.Glamzdin ROCs and DAQ screens for run 9681
  • 344557 12/15/10 16:58 O.Glamazdin Scalers for run 9681
  • 344556 12/15/10 16:10 Aniol hall A tools
  • 344553 12/15/10 16:05 O.Glamazdin Scalers for run 9680
  • 344552 12/15/10 16:04 O.Glamazdin ROCs and DAQ screens for run 9680
  • 344551 12/15/10 16:03 O.Glamazdin ROCs and DAQ screens for run 9680
  • 344549 12/15/10 16:00 Hisham Day shift summary
  • 344548 12/15/10 15:56 yez target screen snapshots
  • 344547 12/15/10 15:09 Hisham HV trip during run # 9679
  • 344542 12/15/10 14:41 garibaldi scalers run 9678
  • 344541 12/15/10 14:40 garibaldi rocs run 9678
  • 344540 12/15/10 14:16 yez target screen snapshots
  • 344537 12/15/10 13:27 garibaldi scalers run 9677
  • 344536 12/15/10 13:26 garibaldi rocs run 9677
  • 344533 12/15/10 12:32 garibaldi scalers run 9676
  • 344532 12/15/10 12:31 garibaldi rocs run 9676
  • 344531 12/15/10 12:16 hisham DVCS DC levels
  • 344530 12/15/10 12:15 yez target screen snapshots
  • 344529 12/15/10 11:57 M Friend Compton disk full
  • 344521 12/15/10 11:34 garibaldi scalers run 9675
  • 344520 12/15/10 11:33 garibaldi rocs run 9675
  • 344517 12/15/10 10:39 garibaldi scalers run 9674
  • 344516 12/15/10 10:38 garibaldi rocs run 9674
  • 344515 12/15/10 10:20 yez Target Screen: presures of both Loop1 and Loop2 go a little high
  • 344513 12/15/10 09:27 garibaldi ROCs run # 9673
  • 344511 12/15/10 09:26 garibaldi scalers run 9673
  • 344508 12/15/10 08:09 G.M. Urciuol 12/15/2010 owl shift report
  • 344501 12/15/10 06:13 siqbal target status
  • 344500 12/15/10 06:00 A. Camsonne Trigger in production
  • 344499 12/15/10 05:52 Carlos Golden run updated to 9670
  • 344498 12/15/10 05:48 G.M. Urciuol large charge asymmetry run
  • 344497 12/15/10 05:38 G. M. Urciuo Calorimeter ADC readback
  • 344496 12/15/10 05:33 A. Camsonne Datamon update attempt
  • 344495 12/15/10 05:28 A. Camsonne Scaler rates with threshold of 80
  • 344494 12/15/10 05:25 A. Camsonne Calorimeter trigger pedestal file updated to calo_ped_9666.txt starting run 9670
  • 344487 12/15/10 04:59 A. Camsonne Run 9666 calorimeter trigger pedestal run 9666
  • 344483 12/15/10 04:53 A. Camsonne Trigger pedestal file updated for gatewidth of 25
  • 344477 12/15/10 04:24 siqbal target status
  • 344464 12/15/10 02:36 siqbal target status
  • 344452 12/15/10 01:14 itard L-arm S1 L-PMT TDC (0-5) peak shifted
  • 344449 12/15/10 00:25 A. Camsonne ND set to 0xC
  • 344446 12/15/10 00:22 A. Camsonne ARS ND set to 0xf
  • 344444 12/15/10 00:00 Holmstrom Swing Shift Summary
  • Tuesday

  • 344443 12/14/10 23:59 Aniol target charts
  • 344441 12/14/10 23:50 A. Camsonne Threshold set to 0 for run 9650
  • 344438 12/14/10 23:25 O.Glamazdin TS11 crush
  • 344437 12/14/10 23:21 A. Camsonne ARS ND set to 8 for trigger timing study
  • 344436 12/14/10 23:19 A. Camsonne Additionnal scaler for dead time measurement
  • 344433 12/14/10 23:04 O.Glamazdin ROCs and DAQ screens for run 9648
  • 344432 12/14/10 23:04 O.Glamazdin ROCs and DAQ screens for run 9648
  • 344431 12/14/10 22:38 O.Glamazdin Scalers for run 9648
  • 344429 12/14/10 21:54 O.Glamazdin Run 9646 was ended due to T11 crush
  • 344427 12/14/10 21:48 Carlos I meant IN
  • 344424 12/14/10 21:15 folts/tech polarity change
  • 344423 12/14/10 21:13 V. Sulkosky/ Calorimeter to Kin 2-High Energy
  • 344422 12/14/10 21:06 Carlos Left collimator out
  • 344420 12/14/10 20:44 Aniol target - change to LD2, Loop1
  • 344418 12/14/10 19:56 O.Glamazdin Scalers for run 9644
  • 344417 12/14/10 19:55 O.Glamazdin ROCs and DAQ screens for run 9644
  • 344416 12/14/10 19:54 O.Glamazdin ROCs and DAQ screens for run 9644
  • 344415 12/14/10 19:45 Aniol target charts
  • 344407 12/14/10 19:10 Holmstrom Hall A Tools
  • 344406 12/14/10 19:03 O.Glamazdin Scalers for run 9643
  • 344405 12/14/10 19:02 Holmstrom Hall A Tools
  • 344404 12/14/10 18:59 O.Glamazdin ROCs and DAQ screens for run 9643
  • 344403 12/14/10 18:59 O.Glamazdin ROCs and DAQ screens for run 9643
  • 344402 12/14/10 18:53 efuchey taking back Compton data
  • 344399 12/14/10 18:09 O.Glamazdin ROCs and DAQ screens for run 9642
  • 344398 12/14/10 18:08 O.Glamazdin ROCs and DAQ screens for run 9642
  • 344397 12/14/10 18:08 O .Glamazdin Scalers for run 9642
  • 344396 12/14/10 18:07 efuchey Stopping compton for the moment
  • 344393 12/14/10 17:19 O.Glamazdin Run 9641 was ended due to DVCS IOC crush
  • 344392 12/14/10 17:12 O.Glamazdin ROCs and DAQ screens for run 9641
  • 344391 12/14/10 17:11 O.Glamazdin ROCs and DAQ screens for run 9641
  • 344390 12/14/10 17:10 O.Glamazdin Scalers for run 9641
  • 344389 12/14/10 16:47 C. Munoz/J. Day Shift Summary (Late Entry)
  • 344388 12/14/10 16:43 Carlos Run 9641: threshold changes
  • 344384 12/14/10 16:06 O.Glamazdin Scalers for run 9640
  • 344383 12/14/10 16:05 O.Glamazdin ROCs and DAQ screens for run 9640
  • 344382 12/14/10 16:04 O.Glamazdin ROCs and DAQ screens for run 9640
  • 344381 12/14/10 16:03 jroche settings for elastic calibration setting 2
  • 344380 12/14/10 15:59 jroche Calo DC currents
  • 344375 12/14/10 14:08 X. Zheng Target status
  • 344370 12/14/10 13:15 X. Zheng Target moved to loop 2 LH2 15 cm
  • 344369 12/14/10 12:55 A. Camsonne Calorimeter pedestal file updated
  • 344364 12/14/10 12:36 A. Camsonne Changing trigger threshold for the calorimeter
  • 344363 12/14/10 12:31 A. Camsonne ARS valid sent to ARS - New firmware testing
  • 344362 12/14/10 12:22 A. Camsonne Trigger files changed
  • 344361 12/14/10 12:17 A. Camsonne Trigger files changed
  • 344358 12/14/10 12:06 atrig Trigger-Download
  • 344357 12/14/10 12:05 atrig Trigger-Download
  • 344356 12/14/10 12:04 A. Camsonne S1 TDC cable swapped
  • 344351 12/14/10 11:45 X. Zheng Target fan speed Loop2 changed to 19%
  • 344339 12/14/10 11:27 jroche LHRS to 2.406 GeV, postive polarity
  • 344337 12/14/10 10:56 hsmith Left Dipole polarity change
  • 344329 12/14/10 10:31 jroche, muno calorimeter to 22.6 deg and 5.5 m
  • 344325 12/14/10 10:24 X. Zheng Target status
  • 344322 12/14/10 10:15 jroche LHRS to 37 deg
  • 344320 12/14/10 10:05 jpchen One pressure read-out unstable
  • 344316 12/14/10 09:52 R. Michaels update to DAQ alarm
  • 344315 12/14/10 09:52 A. Camsonne Writing to disk disabled for TrigDVCS
  • 344311 12/14/10 09:35 A. Camsonne Dec 13 version 3 trigger firmware loaded run 9616
  • 344310 12/14/10 09:24 Carlos Left collimator out
  • 344309 12/14/10 09:16 Carlos Hall is in CONTROLLED ACCESS
  • 344308 12/14/10 08:50 X. Zheng Target fan speed 10% heater 30W
  • 344307 12/14/10 08:38 X. Zheng Target moved to empty
  • 344306 12/14/10 08:36 Hisham Owl shift summary
  • 344305 12/14/10 08:34 Carlos L-HRS magnets going down
  • 344303 12/14/10 08:27 hisham scalers run # 9615
  • 344302 12/14/10 08:27 Hisham ROCs run#9615
  • 344301 12/14/10 08:25 X. Zheng Target status
  • 344300 12/14/10 08:15 X. Zheng Target alarm -- haH2_P_Ballast_R
  • 344296 12/14/10 07:14 garibaldi scalers run 9614
  • 344295 12/14/10 07:13 garibaldi rocs run 9614
  • 344292 12/14/10 06:15 garibaldi scalers run 9613
  • 344291 12/14/10 06:14 garibaldiroc rocs run 9613
  • 344290 12/14/10 05:58 siqbal target status
  • 344288 12/14/10 05:17 hisham DVCS DC levels
  • 344287 12/14/10 05:16 garibaldi scalers run 9612
  • 344286 12/14/10 05:14 garibaldi rocs run 9612
  • 344285 12/14/10 04:42 Hisham Coda crashed during run 9611
  • 344284 12/14/10 04:23 adaq scalers run#9611
  • 344282 12/14/10 04:16 garibaldi rocs run 9611
  • 344281 12/14/10 04:04 siqbal target status
  • 344273 12/14/10 03:22 garibaldi Scalers for run 9610
  • 344272 12/14/10 03:21 garibaldi rocs run n. 9610
  • 344268 12/14/10 02:19 garibaldi scalers run n. 9609
  • 344267 12/14/10 02:18 garibaldi rocs run n. 9609
  • 344266 12/14/10 01:57 siqbal target status
  • 344265 12/14/10 01:31 A. Camsonne Snapshot telnet DAQ lab
  • 344264 12/14/10 01:18 A. Camsonne ARS trigger valid set to 0x2d
  • 344260 12/14/10 01:13 A. Camsonne Production trigger firmware loaded Run 9607
  • 344259 12/14/10 01:12 garibaldi Scalers for run #9607
  • 344257 12/14/10 01:10 garibaldi rocs run n. 9607
  • 344253 12/14/10 00:01 Aniol target charts
  • 344251 12/14/10 00:00 Yi Qiang Alex is changing configuration again :S
  • Monday

  • 344250 12/13/10 23:52 Yi Qiang Swing Shift Summary
  • 344247 12/13/10 23:12 jroche data confirm that the downstream BCM loss 5% in gain
  • 344246 12/13/10 23:05 O.Glamazdin Scalers for run 9604
  • 344245 12/13/10 23:04 O.Glamazdin ROCs and DAQ screens for run 9604
  • 344244 12/13/10 23:04 O.Glamazdin ROCs and DAQ screens for run 9604
  • 344243 12/13/10 23:01 A. Camsonne Scalers with threshold at 0 Run 9604
  • 344240 12/13/10 22:55 A. Camsonne Empty FIFOs in Run 9603 without buffering
  • 344239 12/13/10 22:51 A. Camsonne Scalers run 9603 - 3u A - no buffering
  • 344236 12/13/10 22:48 A. Camsonne Scaler rates with coincidence trigger
  • 344235 12/13/10 22:45 A. Camsonne Dead time estimate at 3uA
  • 344234 12/13/10 22:42 A. Camsonne Empty FIFOs run 9602 ARS valid set to 0x2b
  • 344230 12/13/10 22:22 Magali stait of signal : firmware version 3 of today
  • 344226 12/13/10 22:18 A. Camsonne Empty FIFOs with version 3
  • 344224 12/13/10 22:13 A. Camsonne Test Trigger Firmware December 13 version 3 Run 9599
  • 344223 12/13/10 22:12 Yi Qiang Alex starts the DAQ test again
  • 344221 12/13/10 21:24 O.Glamazdin Scalers for run 9598.
  • 344220 12/13/10 21:24 O.Glamazdin ROCs and DAQ screens for run 9598
  • 344219 12/13/10 21:23 O.Glamazdin ROCs and DAQ screens for run 9598
  • 344217 12/13/10 21:14 Yi Qiang Asked 2 uA beam for run 9598
  • 344214 12/13/10 20:48 Aniol target-increased JT on L2
  • 344213 12/13/10 20:26 O.Glamazdin Scalers for run 9597
  • 344212 12/13/10 20:26 O.Glamazdin ROCs and DAQ screens for run 9597
  • 344211 12/13/10 20:25 O.Glamazdin ROCs and DAQ screens for run 9597
  • 344210 12/13/10 20:14 Yi Qiang DVCS IOC Rebooted
  • 344207 12/13/10 20:02 O.Glamazdin Run 9596 is ended due to DVCS IOC crush
  • 344206 12/13/10 19:54 O.Glamazdin Scalers for Run 9596
  • 344205 12/13/10 19:54 O.Glamazdin ROCs and DAQ screens for Run 9596
  • 344204 12/13/10 19:53 O.Glamazdin ROCs and DAC screens for Run 9596
  • 344196 12/13/10 18:57 O.Glamazdin Scalers for Run 9595
  • 344195 12/13/10 18:56 O.Glamazdin ROCs and DAQ screens for RUN 9595
  • 344194 12/13/10 18:56 O.Glamazdin ROCs and DAQ screens for Run 9595
  • 344191 12/13/10 17:55 O.Glamazdin Scalers for Run 9594
  • 344190 12/13/10 17:54 O.Glamazdin ROCs and DAQ screens for Run 9594
  • 344189 12/13/10 17:53 O.Glamazdin ROCs and DAC for Run 9594
  • 344188 12/13/10 17:51 A. Camsonne Reloaded trigger production firmware
  • 344186 12/13/10 17:49 Yi Qiang Back to production with 1.5 uA
  • 344182 12/13/10 17:41 Yi Qiang Beam current -> 2 uA
  • 344179 12/13/10 17:26 A. Camsonne Loaded slower version od Dec13 trigger firmware Run 9591
  • 344176 12/13/10 17:21 A. Camsonne Run with calorimeter trigger 9590
  • 344175 12/13/10 17:20 Yi Qiang Production paused for DAQ test
  • 344173 12/13/10 17:14 A. Camsonne December 13 Trigger firmware loaded
  • 344171 12/13/10 16:45 O.Glamazdin Scalers for Run 9589
  • 344170 12/13/10 16:45 O.Glamazdin ROCs and DAQ for Run 9589
  • 344169 12/13/10 16:44 O.Glamazdin ROCs and DAC for Run 9589
  • 344168 12/13/10 16:30 R. Michaels new DAQ alarm
  • 344164 12/13/10 16:00 Holmstrom Day shift summary
  • 344163 12/13/10 15:56 O.Glamazdin Scalers screen for Run 9588
  • 344162 12/13/10 15:54 O.Glamazdin ROCs and Run Control screens for Run 9588
  • 344161 12/13/10 15:52 O.Glamazdin Rocs & Runcontrol of a run 9588
  • 344160 12/13/10 15:52 jpchen cryotarget status.
  • 344158 12/13/10 15:27 elfassi Scalers of a run 9587
  • 344157 12/13/10 15:27 elfassi Rocs & Runcontrol of a run 9587
  • 344154 12/13/10 14:27 adaq Scalers for a run 9586
  • 344153 12/13/10 14:27 elfassi Rocs & Runcontrol of a run 9586
  • 344152 12/13/10 14:16 elfassi Dead time for current run 9586
  • 344151 12/13/10 13:55 elfassi BCM rates distribution after the BCM calibration run
  • 344150 12/13/10 13:51 elfassi LHRS Pion Rejector ADC plot
  • 344147 12/13/10 13:20 jroche BCM checks
  • 344145 12/13/10 12:08 jroche followup on the bcm gain jump
  • 344139 12/13/10 11:31 elfassi Rocs & Runcontrol of a run 9584
  • 344137 12/13/10 11:07 elfassi Rocs & Runcontrol of a run 9583
  • 344134 12/13/10 10:34 elfassi Rocs & Runcontrol of a run 9582
  • 344131 12/13/10 10:25 elfassi Scalers of a run 9581
  • 344130 12/13/10 10:25 elfassi Rocs & Runcontrol of a run 9581
  • 344126 12/13/10 09:16 elfassi Scalers of a run 9580
  • 344125 12/13/10 09:15 elfassi Rocs & Runcontrol of a run 9580
  • 344124 12/13/10 08:38 elfassi Wrong scalers timing of a run 9579!
  • 344123 12/13/10 08:30 jpchen Loop 2 fan controller
  • 344122 12/13/10 08:23 jpchen Changed LH2 fan from 24Hz to 46 Hz and LD2 from 60 to 34 Hz.
  • 344120 12/13/10 08:17 elfassi Scalers of run 9579
  • 344119 12/13/10 08:16 elfassi Rocs & Runcontrol for a run 9579
  • 344118 12/13/10 08:01 Hisham Owl shift summary
  • 344114 12/13/10 07:21 de leo run 9578 scalers
  • 344113 12/13/10 07:20 de leo run 9578 rocs
  • 344112 12/13/10 07:00 siqbal target status
  • 344111 12/13/10 06:20 Hisham DVCS IOC rebooted
  • 344108 12/13/10 06:05 hisham DVCS DC levels
  • 344105 12/13/10 05:48 siqbal target moved to hydrogen
  • 344103 12/13/10 05:41 de leo run 9575 scalers
  • 344102 12/13/10 05:40 de leo run 9575 rocs
  • 344099 12/13/10 05:17 de leo run 9574 scalers
  • 344098 12/13/10 05:16 de leo run 9574 rocs
  • 344097 12/13/10 05:01 siqbal target status
  • 344094 12/13/10 04:26 de leo run 9573 scalers
  • 344093 12/13/10 04:25 de leo run 9573 rocs
  • 344085 12/13/10 03:29 de leo run 9572 scalers
  • 344084 12/13/10 03:28 de leo run 9572 rocs
  • 344083 12/13/10 03:14 siqbal target status
  • 344081 12/13/10 02:12 de leo run 9569 scalers
  • 344080 12/13/10 02:11 de leo run 9569 rocs
  • 344079 12/13/10 01:49 Hisham Compton HV off
  • 344077 12/13/10 01:01 de leo run 9568 scalers
  • 344076 12/13/10 01:00 deleo run 9568 rocs
  • 344072 12/13/10 00:01 LeRose Shift Summary (swing)
  • Sunday

  • 344071 12/12/10 23:57 hyao target snapshot
  • 344070 12/12/10 23:38 LeRose DVCS IOC reboot
  • 344065 12/12/10 22:53 garibaldi scalers run n. 9565
  • 344064 12/12/10 22:51 garibaldi rocs run. 9565
  • 344063 12/12/10 22:06 hyao compton high rates
  • 344062 12/12/10 21:55 hyao target snapshot
  • 344056 12/12/10 21:01 garibaldi scalers run n. 9563
  • 344055 12/12/10 21:00 garibaldi rocs 9563
  • 344052 12/12/10 19:53 garibaldi scalrs run 9652
  • 344051 12/12/10 19:52 garibaldi rocs run 9652
  • 344045 12/12/10 19:29 A. Camsonne 7m of delay added on Gas Cernkov - Left arm gas Cerenkov amplitude close to old reference run
  • 344044 12/12/10 19:24 hyao target snapshot
  • 344043 12/12/10 19:15 A. Camsonne Missing S1 TDC
  • 344042 12/12/10 19:15 LeRose Calorimeter Info
  • 344040 12/12/10 18:55 hyao target to L1 15cm
  • 344036 12/12/10 18:49 A. Camsonne Delay setup
  • 344027 12/12/10 18:24 Hisham Delays on Triggers
  • 344019 12/12/10 17:16 hyao target moved to empty
  • 344017 12/12/10 16:13 jroche Another jump in the reading of the downstream BCM ?
  • 344015 12/12/10 15:58 elfassi Target's GUIs & strip charts at the end of day shift
  • 344014 12/12/10 15:50 Huber Day shift summary
  • 344013 12/12/10 15:44 Huber Compton log for #24482
  • 344010 12/12/10 15:23 siqbal #9549 Scalers
  • 344009 12/12/10 15:22 siqbal #9549 RunControl and ROCs
  • 344008 12/12/10 14:37 elfassi Target's GUIs and strip charts @ 2:35PM
  • 344005 12/12/10 14:23 siqbal #9548 Scalers
  • 344004 12/12/10 14:22 siqbal #9548 ROCS and RunControl
  • 344003 12/12/10 14:12 Huber Calo Anode Readback run# 9548
  • 344000 12/12/10 13:23 siqbal #9547 scalers
  • 343999 12/12/10 13:21 siqbal #9547 RunControl and ROCS
  • 343998 12/12/10 12:32 elfassi hal2_Cernox_2_R Fluctuation!
  • 343996 12/12/10 12:19 siqbaq #9546 scalers
  • 343995 12/12/10 12:18 siqbal #9546 RunControl and ROCS
  • 343993 12/12/10 12:04 Huber Hall C bleedthrough
  • 343992 12/12/10 12:02 elfassi Target's GUIs and strip charts @ Noon
  • 343991 12/12/10 11:50 Huber Turn Compton laser off
  • 343985 12/12/10 11:34 Huber Compton rates now very high
  • 343982 12/12/10 11:24 siqbal #9545 scalers
  • 343981 12/12/10 11:24 siqbal #9545 ROCS and RunControl
  • 343980 12/12/10 11:00 Huber Compton stripchart during #24474
  • 343977 12/12/10 10:20 siqbal #9544 Scalers
  • 343976 12/12/10 10:20 siqbal #9544 ROCS and RunControl
  • 343975 12/12/10 09:41 Huber Compton run# 24471 stripchart
  • 343972 12/12/10 09:20 iqbal #9543 scalars
  • 343971 12/12/10 09:20 siqbal #9543 ROCS and RunControl
  • 343970 12/12/10 09:10 Huber FB ON - beam positions now more stable
  • 343969 12/12/10 09:01 elfassi Target's strip charts
  • 343968 12/12/10 08:56 elfassi Target Main GUIs
  • 343967 12/12/10 08:53 Huber Fast Feedback Status
  • 343966 12/12/10 08:36 siqbal #9542 - Lumi4 problem
  • 343965 12/12/10 08:33 siqbal run #9542 - problem pion rejector ADC
  • 343964 12/12/10 08:25 Huber Compton - variable rates
  • 343961 12/12/10 08:17 siqbal Run #9542: Scalers
  • 343960 12/12/10 08:16 siqbal Run #9542: ROCS and RunControl
  • 343958 12/12/10 08:00 igorko shift summary
  • 343957 12/12/10 07:25 igorko compton info
  • 343956 12/12/10 07:23 igorko calorimeter info
  • 343953 12/12/10 07:14 de leo run 9541 scalers
  • 343952 12/12/10 07:13 de leo run 9541 rocs
  • 343951 12/12/10 07:07 igorko target info
  • 343948 12/12/10 06:08 de leo run 9540 scalers
  • 343947 12/12/10 06:07 de leo run 9540 rocs
  • 343944 12/12/10 05:04 de leo run 9539 scalers
  • 343943 12/12/10 05:03 igorko target info
  • 343942 12/12/10 05:03 de leo run 9539 rocs
  • 343939 12/12/10 03:58 de leo run 9538 scalers
  • 343938 12/12/10 03:57 de leo run 9539 rocs
  • 343932 12/12/10 03:15 Carlos Golden run updated to 9538
  • 343931 12/12/10 03:09 igorko target info
  • 343928 12/12/10 01:58 igorko target movement
  • 343926 12/12/10 01:46 de leo run 9536 scalers
  • 343925 12/12/10 01:45 de leo run 9536 rocs
  • 343924 12/12/10 01:04 igorko target info
  • 343923 12/12/10 00:46 igorko IOC reboot
  • 343919 12/12/10 00:37 igorko target movement
  • 343917 12/12/10 00:36 de leo run 9534 scalers
  • 343916 12/12/10 00:35 de leo run 9534 rocs
  • 343915 12/12/10 00:08 elfassi Target @ Optics position
  • 343912 12/12/10 00:02 hyao Swing Shift Summary
  • Saturday

  • 343908 12/11/10 23:41 elfassi Target's guis & strip charts @ 11:41PM
  • 343906 12/11/10 23:09 garibaldi scalers 9531
  • 343905 12/11/10 23:08 garibaldi rocs 9531
  • 343902 12/11/10 22:05 garibaldi scalers run 9530
  • 343901 12/11/10 22:04 garibaldi rocs run 9530
  • 343900 12/11/10 21:51 Carlos Golden run updated to 9529
  • 343899 12/11/10 21:47 Carlos HRS after timing changes
  • 343896 12/11/10 21:05 garibaldi scalers
  • 343895 12/11/10 21:04 Garibaldi Rocs
  • 343894 12/11/10 21:01 elfassi Loop 2 Fan controls issue
  • 343893 12/11/10 20:25 efuchey Going back to compton production
  • 343892 12/11/10 20:22 efuchey Compton rates
  • 343890 12/11/10 20:05 hyao Beam recovery done and start production
  • 343889 12/11/10 20:05 elfassi Target @ LH2 position
  • 343883 12/11/10 19:31 elfassi Target @ empty position
  • 343881 12/11/10 19:23 A. Camsonne DAQ work
  • 343880 12/11/10 19:20 hyao Trigger work done. Ask for beam permit
  • 343872 12/11/10 18:36 A. Camsonne Timing 1875
  • 343863 12/11/10 17:58 M Friend Compton LED runlist
  • 343850 12/11/10 17:34 M Friend Compton DAQ back to standard config
  • 343840 12/11/10 16:13 elfassi Target's guis and strip charts
  • 343838 12/11/10 16:03 Huber Day shift summary
  • 343835 12/11/10 15:50 M Friend Compton disk full
  • 343834 12/11/10 15:37 M Friend Compton DAQ setup for LED runs
  • 343826 12/11/10 14:16 eschulte Day Shift Target Screens: Main Screen 14:00
  • 343825 12/11/10 14:15 eschulte Day Shift Target Screens: Strip Charts
  • 343805 12/11/10 12:47 A. Camsonne DVCS HRS trigger work / checks
  • 343800 12/11/10 12:09 eschulte Day Shift Target Strip Charts: 12:00
  • 343799 12/11/10 12:08 eschulte Day Shift Target Screens: 12:00
  • 343778 12/11/10 10:12 A. Camsonne ROC17 scaler assignement
  • 343777 12/11/10 10:06 eschulte Day Shift Target Main Screen: 10:00
  • 343776 12/11/10 10:05 eschulte Day Shift Target Strip Charts: 10:00
  • 343775 12/11/10 09:59 A. Camsonne Signal at trigger
  • 343772 12/11/10 08:26 eschulte Day Shift Target Strip Charts: 08:30
  • 343771 12/11/10 08:25 eschulte Day Shift Target Target Screen: 08:30
  • 343763 12/11/10 08:00 igorko Shift summary
  • 343761 12/11/10 07:50 igorko Calorimeter Readback
  • 343750 12/11/10 06:57 A. Camsonne Added ARS configuratution into CODA file as event 137
  • 343744 12/11/10 05:08 n. Muangma target computer reboot
  • 343743 12/11/10 05:07 igorko Beam permit
  • 343741 12/11/10 04:55 A. Camsonne halla-rtr power cycled
  • 343738 12/11/10 04:28 igorko controll access
  • 343731 12/11/10 03:33 A. Camsonne Blaster tests HRS ROCs
  • 343730 12/11/10 03:28 A. Camsonne DAQ very slow ... network issue ?
  • 343722 12/11/10 02:34 n. muangma target position
  • 343711 12/11/10 01:14 Bob/Alex/His trigger timing checks
  • Friday

  • 343708 12/10/10 23:58 Huber Swing shift summary
  • 343707 12/10/10 23:41 A. Camsonne Block 8 had 150 cm delay cable
  • 343706 12/10/10 23:16 Huber Calorimeter Anode Currents
  • 343701 12/10/10 22:05 A. Camsonne Trigger gates width and timing
  • 343700 12/10/10 22:04 A. Camsonne Trigger gates width and timing
  • 343699 12/10/10 22:03 A. Camsonne Trigger gates width and timing
  • 343698 12/10/10 21:37 A. Camsonne Trigger gates cable numbering
  • 343697 12/10/10 21:37 Huber Hall A in Controlled Access
  • 343695 12/10/10 21:21 A. Camsonne Trigger firmware version December 10th loaded
  • 343693 12/10/10 21:01 A. Camsonne Current ARS - HRS timing
  • 343690 12/10/10 20:30 pzhu scaler9456
  • 343689 12/10/10 20:29 pzhu coda9456
  • 343688 12/10/10 19:56 Aniol target charts
  • 343686 12/10/10 19:37 pzhu scaler9455
  • 343684 12/10/10 19:37 pzhu coda9455
  • 343676 12/10/10 18:29 pzhu scaler9454
  • 343675 12/10/10 18:28 pzhu coda9454
  • 343672 12/10/10 17:30 pzhu coda9453
  • 343671 12/10/10 17:30 pzhu scaler9453
  • 343668 12/10/10 16:24 pzhu scaler9452
  • 343667 12/10/10 16:23 Aniol target charts
  • 343666 12/10/10 16:23 pzhu coda9452
  • 343663 12/10/10 15:53 eschulte 10 December 2010 Day Shift Run List
  • 343662 12/10/10 15:53 eschulte Day Shift Log 10 December 2010
  • 343659 12/10/10 15:24 de leo run 9451 scalers
  • 343658 12/10/10 15:23 de leo run 9451 roc
  • 343657 12/10/10 14:54 eschulte Day Shift Target 15:00
  • 343656 12/10/10 14:53 eschulte Day Shift Target Strip Charts 15:00
  • 343653 12/10/10 14:26 de leo run 9450 scalers
  • 343652 12/10/10 14:25 de leo run 9450 roc
  • 343651 12/10/10 14:15 jroche Mott results from yesterday
  • 343650 12/10/10 13:47 eschulte Compton Strip Chart 1
  • 343649 12/10/10 13:42 eschulte Shift Checklist: Calorimeter Currents
  • 343648 12/10/10 13:35 eschulte BPM after recovery
  • 343646 12/10/10 13:25 eschulte Moving target to LH2
  • 343645 12/10/10 13:00 eschulte Back to Compton!
  • 343644 12/10/10 12:57 A. Camsonne Compton rates slightly improved
  • 343643 12/10/10 12:27 eschulte Empty Target
  • 343642 12/10/10 12:22 eschulte MCC to deliver beam
  • 343641 12/10/10 12:21 eschulte Move target to EMPTY: Beam Recovery
  • 343640 12/10/10 12:21 eschulte Day Shift Target Chart 08:00
  • 343639 12/10/10 12:12 eschulte Day Shift Target Screen 12:00
  • 343638 12/10/10 12:12 eschulte Day Shift Target Charts 12:00
  • 343637 12/10/10 11:59 eschulte MCC: Finished with ARC6 now path lengths
  • 343631 12/10/10 11:19 eschulte DVCS Calo IOC Reboot
  • 343630 12/10/10 09:59 eschulte Day Shift Target Pictures 10:00
  • 343629 12/10/10 09:58 eschulte Day Shift Target Graphs 10:00
  • 343626 12/10/10 09:37 de leo run 9448 scalers
  • 343625 12/10/10 09:36 de leo run 9448 roc
  • 343622 12/10/10 08:35 de leo run 9447 scalers
  • 343621 12/10/10 08:34 de leo run 9447 roc
  • 343620 12/10/10 08:08 eschulte Beam Back!
  • 343619 12/10/10 08:05 eschulte Day Shift Target Strip Charts 08:00
  • 343618 12/10/10 08:04 n. muangma owl shift summary
  • 343615 12/10/10 07:27 igorko end of run #9446
  • 343614 12/10/10 07:26 igorko end of the run #9446 info
  • 343610 12/10/10 06:27 igorko end of run#9445 info
  • 343609 12/10/10 06:26 igorko end of run#9445 info
  • 343608 12/10/10 05:56 n muangma target strip chart
  • 343605 12/10/10 05:28 igorko end of run# 9444
  • 343604 12/10/10 05:27 igorko end of run #9444
  • 343603 12/10/10 05:16 igorko ROCs sync check run number #9444
  • 343600 12/10/10 04:24 igorko end of run #9443
  • 343599 12/10/10 04:23 igorko end of run #9443
  • 343598 12/10/10 04:01 igorko run #9443 ROCs sync
  • 343590 12/10/10 03:28 igorko end of run #9442
  • 343589 12/10/10 03:27 igorko end of run #9442
  • 343588 12/10/10 02:50 igorko ROCs Sync check run #9442
  • 343585 12/10/10 02:25 igorko end of run # 9441
  • 343584 12/10/10 02:24 igorko end of run #9441 info
  • 343583 12/10/10 02:08 n.muangma target strip charts
  • 343582 12/10/10 01:47 igorko ROC 14 Sync check
  • 343581 12/10/10 01:42 N Muangma compton HV OFF
  • 343580 12/10/10 01:41 Hisham High rates for Compton ~>4kHz
  • 343577 12/10/10 01:18 igorko end of run 9440
  • 343576 12/10/10 01:17 igorko end of run #9440
  • 343575 12/10/10 01:12 Hisham Compton rates run # 24424
  • 343574 12/10/10 00:44 Carlos Golden run updated to 9440
  • 343571 12/10/10 00:09 igorko end of run #9439
  • 343570 12/10/10 00:09 igorko end of run #9439 scaler info
  • Thursday

  • 343568 12/09/10 23:58 Huber Swing shift summary
  • 343567 12/09/10 23:24 A. Camsonne Reduced pedestal subtraction for Gas Cerenkov for pedestal check
  • 343564 12/09/10 23:04 A. Camsonne Compton rates improved - taking Compton data tonight
  • 343562 12/09/10 22:55 de leo run 9437 scalers
  • 343561 12/09/10 22:54 de leo run 9437 roc
  • 343560 12/09/10 22:34 A. Camsonne Asked to MCC to look at Compton rates
  • 343555 12/09/10 21:30 de leo run 9435 scalers
  • 343554 12/09/10 21:29 de leo run 9435 roc
  • 343553 12/09/10 21:28 A. Camsonne Double peak in the S2m TDCs
  • 343552 12/09/10 21:28 Huber Calo Anode currents for run 9435
  • 343549 12/09/10 20:32 de leo run 9434 scalers
  • 343548 12/09/10 20:31 A. Camsonne Timing of trigger changed
  • 343547 12/09/10 20:31 de leo run 9434 roc
  • 343546 12/09/10 20:29 A. Camsonne Clock trigger set to 11 Hz
  • 343538 12/09/10 19:31 de leo run 9433 scalers
  • 343537 12/09/10 19:15 Aniol target-cernox2 gives unphysical result
  • 343536 12/09/10 19:10 Aniol peculiar L2 temp chart
  • 343535 12/09/10 18:58 de leo some differences run 9433
  • 343532 12/09/10 18:29 A. Camsonne Production firmware loaded for trigger
  • 343528 12/09/10 17:21 efuchey Compton rates for 12-9-2010
  • 343517 12/09/10 16:43 A. Camsonne ND set back to 0
  • 343511 12/09/10 16:00 A. Deur Day shift summary
  • 343504 12/09/10 15:44 A. Deur Snapshot of calorimter anode currents
  • 343497 12/09/10 14:29 A. Deur Target snapshot
  • 343496 12/09/10 14:11 A. Camsonne LED trigger gate time scan Run 9414
  • 343494 12/09/10 13:33 A. Camsonne ARS ND set to 6
  • 343493 12/09/10 13:31 A. Camsonne ARS ND set to 6
  • 343468 12/09/10 11:37 A. Deur Controlled access extended to 1pm
  • 343461 12/09/10 11:13 A. Camsonne Timing at trigger inputs
  • 343460 12/09/10 11:12 A. Camsonne Timing at trigger inputs
  • 343459 12/09/10 10:43 A. Camsonne Gas cerenkov gate width change
  • 343454 12/09/10 08:44 A. Camsonne Calorimeter trigger firmware loaded version of December 8th 2010
  • 343453 12/09/10 08:29 A. Deur Hall is in controlled access
  • 343450 12/09/10 08:01 rakithab Owl shift summary
  • 343446 12/09/10 07:28 rakithab CODA status during the run 9399
  • 343445 12/09/10 07:27 rakithab Hall A scaler data during the run 9399
  • 343442 12/09/10 06:35 rakithab CODA status during the run 9398
  • 343441 12/09/10 06:34 rakithab Hall A scaler data during the run 9398
  • 343440 12/09/10 05:59 VBellini Target snapshots
  • 343437 12/09/10 05:32 rakithab CODA status during the run 9397
  • 343436 12/09/10 05:31 rakithab Hall A scaler data during the run 9397
  • 343433 12/09/10 04:30 rakithab CODA status during the run 9396
  • 343432 12/09/10 04:29 rakithab Hall A scaler data during the run 9396
  • 343431 12/09/10 04:04 VBellini Target snapshots
  • 343423 12/09/10 03:32 VBellini ROCs run # 9395
  • 343422 12/09/10 03:32 rakithab Hall A scaler data during the run 9395
  • 343421 12/09/10 03:03 rakithab Calorimeter anaode current log
  • 343418 12/09/10 02:21 rakithab Hall A scaler data during the run 9394
  • 343417 12/09/10 02:21 rakithab CODA status during the run 9394
  • 343416 12/09/10 02:01 VBellini Target snapshots
  • 343413 12/09/10 01:17 Hisham Runcontrol, ROCs for Run#9393
  • 343412 12/09/10 01:16 hisham Scalers for run #9393
  • 343409 12/09/10 00:47 Hisham Runcontrol, ROCs for run 9392
  • 343408 12/09/10 00:45 hisham Run 9392: scalers
  • 343407 12/09/10 00:16 VBellini Target snapshots
  • Wednesday

  • 343405 12/08/10 23:55 Carlos ROC17 not responding
  • 343404 12/08/10 23:53 Huber Swing shift summary
  • 343401 12/08/10 22:51 Carlos RE: Beam current/target fluctuation
  • 343398 12/08/10 22:48 itard scaler data run 9389
  • 343397 12/08/10 22:48 itard entries end of run 9389
  • 343396 12/08/10 22:11 Huber Turned Compton HV off
  • 343395 12/08/10 21:51 Huber Hall A General Tools
  • 343394 12/08/10 21:40 aniol target-brief excursion to 8uA
  • 343393 12/08/10 21:40 Huber Calorimeter anode current log
  • 343389 12/08/10 21:11 itard entries end of run 9387
  • 343388 12/08/10 21:11 itard scaler data run 9387
  • 343387 12/08/10 20:51 efuchey Rates wild again
  • 343386 12/08/10 20:39 efuchey Go back to Compton production
  • 343385 12/08/10 20:17 efuchey Cavity has been tweaked
  • 343382 12/08/10 20:10 itard entries end of run 9386
  • 343381 12/08/10 20:09 itard scaler data run 9386
  • 343380 12/08/10 19:56 aniol target charts
  • 343379 12/08/10 19:48 efuchey Rates going wild
  • 343371 12/08/10 19:13 itard entries end of run 9385
  • 343370 12/08/10 19:13 itard scaler data run 9385
  • 343369 12/08/10 18:15 itard run 9383 was stopped (problem of disk)
  • 343363 12/08/10 17:50 itard entries end of run 9382
  • 343362 12/08/10 17:50 itard data scaler run 9382
  • 343361 12/08/10 17:30 aniol target graphs
  • 343358 12/08/10 16:51 itard entries end of run 9381
  • 343357 12/08/10 16:51 itard scaler data run 9381
  • 343356 12/08/10 16:31 Carlos Back to /adaql2 only
  • 343355 12/08/10 16:21 Carlos Disks full: writing to adaql1
  • 343353 12/08/10 15:59 hyao Day Shift Summary
  • 343350 12/08/10 15:37 de leo run 9380 scalers
  • 343349 12/08/10 15:36 de leo run 9380 roc
  • 343347 12/08/10 14:40 hyao Beam is back. 1uA
  • 343346 12/08/10 14:35 de leo run 9379 roc
  • 343345 12/08/10 14:19 hyao Elog entry for beam off
  • 343344 12/08/10 14:08 n.muangma target strip charts
  • 343343 12/08/10 14:06 hyao Beam is Off for few minutes
  • 343342 12/08/10 14:04 de leo run 9379 scalers
  • 343338 12/08/10 13:32 de leo run n. 9378 scalers
  • 343337 12/08/10 13:31 de leo run n. 9378 roc
  • 343334 12/08/10 12:36 de leo run 9377 scalers
  • 343333 12/08/10 12:35 de leo run n. 9377 roc
  • 343325 12/08/10 11:23 de leo run 9376 scalers
  • 343324 12/08/10 11:22 de leo run n. 9376 roc
  • 343321 12/08/10 10:28 de leo run n. 9375 scalers
  • 343320 12/08/10 10:26 de leo run n. 9375 roc
  • 343319 12/08/10 09:42 n. muangma target strip charts
  • 343317 12/08/10 09:21 Hisham Time shift in calo ADC
  • 343316 12/08/10 09:21 de leo run n. 9374 scalers
  • 343315 12/08/10 09:19 de leo run n. 9374 roc
  • 343312 12/08/10 08:19 de leo run n. 9373 scalers
  • 343311 12/08/10 08:18 de leo run n. 9373 roc
  • 343310 12/08/10 08:01 M Friend Owl shift summary
  • 343308 12/08/10 08:00 M Friend Compton runlist updated
  • 343307 12/08/10 07:58 VBellini End of Shift target snapshots
  • 343306 12/08/10 07:22 buddhini Sclaer rates at the begingind of run 9373
  • 343304 12/08/10 07:21 buddhini CODA windows before the end of
  • 343302 12/08/10 06:34 M Friend Calorimeter anode currents
  • 343301 12/08/10 06:21 buddhini Scaler rates at the begining of run 9372
  • 343299 12/08/10 06:20 buddhini CODA windows before the end of run 9371
  • 343297 12/08/10 06:10 VBellini Target snapshots
  • 343296 12/08/10 05:25 M Friend Compton rates
  • 343295 12/08/10 05:23 M Friend Compton photon detector centered
  • 343291 12/08/10 05:09 Carlos Golden run updated to 9368
  • 343289 12/08/10 04:55 M Friend Compton photon detector moved
  • 343288 12/08/10 04:54 M Friend Requested beam through the Compton chicane
  • 343285 12/08/10 04:48 Carlos Rates kin1-high
  • 343282 12/08/10 04:41 A. Camsonne Loading trigger production firmware
  • 343280 12/08/10 04:21 A. Camsonne Empy FIFOs ARS with new firmware - run 9366
  • 343277 12/08/10 04:12 A. Camsonne Timing with new firmware
  • 343275 12/08/10 04:00 M Friend Cosmic paddles HV off for beam
  • 343274 12/08/10 04:00 VBellini Targed moved to LH2 15 cm
  • 343273 12/08/10 03:59 M Friend Beam recovery finished
  • 343271 12/08/10 03:54 M Friend Harp scan result
  • 343270 12/08/10 03:52 A. Camsonne ROC18 counters headers changed
  • 343263 12/08/10 03:08 M Friend Beam recovery starting
  • 343262 12/08/10 02:56 adaq CODA windows before the end of run 9363
  • 343260 12/08/10 02:06 VBellini Target snapshots
  • 343259 12/08/10 01:24 M Friend Accelerator status update
  • Tuesday

  • 343255 12/07/10 23:53 A. Camsonne HV on for cosmics data taking
  • 343254 12/07/10 23:43 A. Camsonne Trigger pedestal generated for trigger
  • 343252 12/07/10 23:38 hyao target snapshot
  • 343243 12/07/10 20:28 hyao target snapshot
  • 343233 12/07/10 18:36 A. Camsonne LED calo scan run 9355
  • 343232 12/07/10 18:27 hyao target screen
  • 343213 12/07/10 17:36 Carlos Preliminary survey results
  • 343205 12/07/10 17:04 R. Michaels Compton ROC6 back to normal
  • 343198 12/07/10 16:38 Carlos Kin1-high settings
  • 343196 12/07/10 16:36 R. Michaels info about network change from B. Hess
  • 343188 12/07/10 15:37 jroche LHRS to 16.37 deg
  • 343176 12/07/10 12:05 A. Camsonne boot parameters changed to load updated trigger library
  • 343172 12/07/10 11:47 A. Camsonne Calorimeter trigger firmware version trig-dec06
  • 343166 12/07/10 11:33 Bob Michaels Network issue between Compton and DVCS DAQ solved
  • 343151 12/07/10 08:30 Hisham Hall C charge Asymmertry test during production run 9315
  • 343150 12/07/10 08:26 A. Camsonne Blaster tests with adaql2 and Compton
  • 343149 12/07/10 08:13 Holmstrom Adjusting fan and JT valve
  • 343147 12/07/10 08:00 M Friend Owl shift summary
  • 343145 12/07/10 07:59 VBellini End of shift target snapshots
  • 343141 12/07/10 07:00 de leo run n. 9315 scaler
  • 343140 12/07/10 06:59 M Friend Beam down for beam studies
  • 343139 12/07/10 06:58 de leo run n. 9315 roc
  • 343138 12/07/10 06:13 Carlos Check disk space
  • 343136 12/07/10 06:04 M Friend DVCS ioc rebooted
  • 343135 12/07/10 05:59 de leo run 9314 stopped
  • 343131 12/07/10 05:54 deleo run 9213 stopped
  • 343130 12/07/10 05:53 Carlos Changed to adaql2
  • 343127 12/07/10 05:49 de leo run n. 9312 scaler
  • 343126 12/07/10 05:49 de leo run n. 9312 roc
  • 343125 12/07/10 05:40 VBellini Target snapshots
  • 343124 12/07/10 05:35 deleo run 9312 and previous
  • 343121 12/07/10 04:52 de leo run n. 9311 scaler
  • 343120 12/07/10 04:51 de leo run n. 9311 roc
  • 343119 12/07/10 04:12 M Friend Calorimeter PMT currents
  • 343116 12/07/10 03:54 de leo run 0310 scaler
  • 343115 12/07/10 03:54 de leo run 9310 roc
  • 343109 12/07/10 03:08 VBellini Target snapshots
  • 343106 12/07/10 02:50 de leo run 9309 scalers
  • 343105 12/07/10 02:49 de leo run 9309 roc
  • 343104 12/07/10 01:52 M Friend Disk full, changed pingpong to USEADAQL1=1
  • 343102 12/07/10 01:48 M Friend end of run 9308 (disk full, coda crashed)
  • 343100 12/07/10 01:12 M Friend Large charge asym during run 9305
  • 343099 12/07/10 01:05 VBellini Target back to L1 D2 15cm
  • 343095 12/07/10 00:50 de leo run n, 9307
  • 343094 12/07/10 00:48 de leo run n. 9307
  • 343092 12/07/10 00:25 VBellini Target moved to optics
  • 343090 12/07/10 00:15 Hisham Runcontrol, prescales, ROCs for run 9306
  • 343089 12/07/10 00:14 hisham Scalers run 9306
  • Monday

  • 343087 12/06/10 23:54 Huber Swing shift summary
  • 343084 12/06/10 23:37 itard entries end of run 9305
  • 343083 12/06/10 23:37 itard sacler data 9305
  • 343082 12/06/10 23:26 Yi Qiang Target screens
  • 343079 12/06/10 22:40 itard scaler data run 9304
  • 343078 12/06/10 22:39 itard entries end of run 9304
  • 343077 12/06/10 21:56 Carlos Please STOP the run BEFORE rebooting the DVCS IOC
  • 343072 12/06/10 20:55 itard scaler data run 9302
  • 343071 12/06/10 20:55 itard entries end of run 9302
  • 343068 12/06/10 19:49 itard scaler data 9301
  • 343067 12/06/10 19:49 itard entries end of run 9301
  • 343058 12/06/10 18:50 itard scaler data run 9300
  • 343057 12/06/10 18:50 itard entries end of run 9300
  • 343054 12/06/10 17:46 itard scaler data run 9299
  • 343053 12/06/10 17:46 itard entries end of run 9299
  • 343052 12/06/10 17:36 pzhu target statement
  • 343050 12/06/10 16:45 itard entries end of run 9298
  • 343048 12/06/10 16:02 jpchen day shift summary
  • 343045 12/06/10 15:36 hyao Remove some unwanted lines
  • 343044 12/06/10 15:28 hyao Cyro charts 2
  • 343043 12/06/10 15:27 hyao Cyro charts
  • 343042 12/06/10 15:27 hyao L2 charts
  • 343041 12/06/10 15:26 hyao L1+L2 charts
  • 343040 12/06/10 15:26 hyao L1 charts
  • 343039 12/06/10 15:24 hyao Fan speed of LH2 15cm changed to 24Hz
  • 343036 12/06/10 14:05 hyao target gui rebooted
  • 343035 12/06/10 13:21 hyao target snapshot
  • 343032 12/06/10 13:14 de leo run n. 9295
  • 343031 12/06/10 13:13 de leo run n. 9295
  • 343028 12/06/10 12:08 de leo run n. 9294
  • 343027 12/06/10 12:07 de leo run n. 9294
  • 343026 12/06/10 11:43 R. Michaels test of deadtime (late entry)
  • 343018 12/06/10 11:01 de leo run 9293
  • 343017 12/06/10 11:00 de leo run 9293
  • 343016 12/06/10 10:57 Carlos Compton DAQ tests: run 9293
  • 343015 12/06/10 10:05 hyao target screen
  • 343014 12/06/10 09:56 Hisham target moved to LD2
  • 343012 12/06/10 09:53 hyao target moved to L1 15cm
  • 343011 12/06/10 09:48 deleo scalers run n. 9292
  • 343009 12/06/10 09:47 de leo run control, rocs run n. 9292
  • 343007 12/06/10 08:39 jpchen Loop 2 fan using loop 3 controller now
  • 343005 12/06/10 07:58 M Friend Owl shift summary
  • 343004 12/06/10 07:13 M Friend DVCS ioc reboot
  • 343003 12/06/10 07:00 jpchen loop 2 stopped working, could not reset.
  • 343002 12/06/10 06:00 VBellini Run stop due to target red alarm
  • 343000 12/06/10 05:57 M Friend Called for controlled access
  • 342999 12/06/10 05:47 VBellini Loop2 in trouble: fan 0 Hz
  • 342996 12/06/10 05:29 VBellini Scalers End of Run #9290
  • 342995 12/06/10 05:27 VBellini ROCs End of Run #9290
  • 342992 12/06/10 04:33 VBellini Scalers End of Run #9289
  • 342991 12/06/10 04:31 VBellini ROCs End of Run #9289
  • 342990 12/06/10 04:13 M Friend Calorimeter current snapshot
  • 342989 12/06/10 04:03 VBellini Target snapshots
  • 342980 12/06/10 03:28 VBellini Scalers End of Run #9288
  • 342979 12/06/10 03:26 VBellini ROCS End of Run #9288
  • 342976 12/06/10 02:28 VBellini Scalers End of Run #9287
  • 342975 12/06/10 02:27 VBellini ROCS End of Run #9287
  • 342974 12/06/10 02:25 Carlos Back to /adaql2 only
  • 342973 12/06/10 02:05 Carlos Writing to adaql1
  • 342972 12/06/10 01:58 VBellini Target screens
  • 342969 12/06/10 01:27 VBellini Scalers End of Run #9286
  • 342968 12/06/10 01:25 VBellini ROCS End of RUN #9286
  • 342967 12/06/10 00:44 M Friend Calorimiter pedestal shift
  • 342964 12/06/10 00:22 hisham Runcontrol, prescales, ROCs for eff run 9285
  • 342963 12/06/10 00:21 hisham scalers for eff. run 9285
  • Sunday

  • 342961 12/05/10 23:56 Ramesh Subed Shift Summary
  • 342958 12/05/10 23:38 hyao target tool Fan=46Hz
  • 342957 12/05/10 23:30 Huber RunControl, ROCs, Prescale for 9284
  • 342956 12/05/10 23:29 Huber Scalers for efficiency run 9284
  • 342953 12/05/10 22:45 Huber Scalers for run 9283
  • 342952 12/05/10 22:44 Huber RunControl, ROCs, prescale for run 9283
  • 342951 12/05/10 22:25 Ramesh Beam back at 22:18
  • 342950 12/05/10 22:23 Ramesh calorimeter IOC tripped
  • 342948 12/05/10 22:05 Ramesh Beam down for 10-15 minutes
  • 342944 12/05/10 21:20 Huber Scalers for run 9281
  • 342943 12/05/10 21:19 Huber RunControl, ROCs, prescales for run 9281
  • 342942 12/05/10 21:18 hyao target tool and temp graph
  • 342941 12/05/10 21:17 hyao target pressure and power graph
  • 342940 12/05/10 21:17 Ramesh Calorimeter PMT current
  • 342939 12/05/10 20:21 hyao oscillation of hal2_Cernox_2_R
  • 342936 12/05/10 20:11 Huber scalers for run 9280
  • 342935 12/05/10 20:11 Huber RunControl, ROCs, prescale for run 9280
  • 342927 12/05/10 19:12 Huber Scalers for run 9279
  • 342926 12/05/10 19:11 Huber RunControl, ROCs, prescales for run 9279
  • 342924 12/05/10 18:17 Huber Scalers for run 9278
  • 342923 12/05/10 18:15 Huber RunControl, ROCs, prescales for run 9278
  • 342920 12/05/10 17:23 Huber Scalers for run 9277
  • 342919 12/05/10 17:22 Huber RunControl, ROCs, prescales for run 9277
  • 342917 12/05/10 16:14 hyao target from L1 15cm to L2 15cm
  • 342913 12/05/10 16:09 Huber Scalers for run 9275
  • 342912 12/05/10 16:08 Huber RunControl, prescales for run 9275
  • 342910 12/05/10 16:04 bogdanw shift summary
  • 342909 12/05/10 15:11 Lindgren roc 17 crash
  • 342908 12/05/10 15:11 igorko target movement
  • 342906 12/05/10 15:08 A. Camsonne ADC gate timing
  • 342904 12/05/10 14:59 Lindgren eor 9274 rocs and coda
  • 342903 12/05/10 14:22 bogdanw calorimeter currents
  • 342901 12/05/10 13:58 igorko target info
  • 342898 12/05/10 12:54 Lindgren Roc 17 not reported reof rocs and coda
  • 342897 12/05/10 12:50 R. Michaels re: timing shifts
  • 342896 12/05/10 12:22 bogdanw HARP is ready for scan
  • 342893 12/05/10 12:11 lindgren eor 9271 scalers
  • 342892 12/05/10 12:09 Lindgren eor 9271 rocs and coda
  • 342891 12/05/10 12:08 bogdanw beam strip s
  • 342890 12/05/10 11:48 R. Lindgren IOC rebooted for calor scalers
  • 342889 12/05/10 11:38 bogdanw mid shift summary
  • 342888 12/05/10 11:33 Carlos PLEASE LOG ACTIVITY GOING ON
  • 342881 12/05/10 11:06 igorko target movement
  • 342880 12/05/10 10:42 igorko target movement
  • 342879 12/05/10 10:29 igorko target info
  • 342878 12/05/10 09:06 Lindgren Rocs and Coda eor 9270
  • 342877 12/05/10 09:04 igorko target frammable gas sensor alarm
  • 342876 12/05/10 09:03 lindgren scalers at eor 9270
  • 342872 12/05/10 08:23 Lindgren End of Run 9269 Scalers
  • 342871 12/05/10 08:22 Lindgren Rocs and coda Run 9269
  • 342869 12/05/10 08:00 VBellini End of run target snapshots
  • 342868 12/05/10 07:57 M Friend Owl shift summary
  • 342867 12/05/10 07:35 jpchen loop 1 back to 22K/23psj
  • 342866 12/05/10 07:27 VBellini Loop1 looks fine now
  • 342863 12/05/10 07:22 hkkang scaler log of 9268
  • 342862 12/05/10 07:21 hkkang end-of-run 9268
  • 342861 12/05/10 06:57 jpchen Loop 1 fan reset to 34 Hz/14%, cooling down now
  • 342860 12/05/10 06:40 VBellini Loop1 screenshot
  • 342859 12/05/10 06:27 VBellini Loop 1 sensors
  • 342856 12/05/10 06:22 hkkang scaler log 9267
  • 342855 12/05/10 06:21 hkkang end-of-run 9267
  • 342854 12/05/10 06:06 M Friend Loop 1 went bad
  • 342853 12/05/10 05:53 VBellini Target screens
  • 342850 12/05/10 05:21 hkkang scaler log of run9266
  • 342849 12/05/10 05:20 hkkang end-of-run 9266
  • 342848 12/05/10 04:30 M Friend Calorimeter currents
  • 342845 12/05/10 04:20 hkkang scaler info 9265
  • 342844 12/05/10 04:19 hkkang end-of-run 9265
  • 342843 12/05/10 04:15 VBellini Target screens
  • 342842 12/05/10 03:41 hkkang BPMA y peak shifted
  • 342834 12/05/10 03:20 hkkang scaler log of 9264
  • 342833 12/05/10 03:19 hkkang end-of-run 9264
  • 342832 12/05/10 02:56 M Friend Fast feedback was off
  • 342830 12/05/10 02:21 M Friend DVCS ioc reboot
  • 342828 12/05/10 02:16 hkkang scaler 9263
  • 342827 12/05/10 02:15 hkkang end-of-run 9263
  • 342826 12/05/10 01:42 VBellini Target screens
  • 342825 12/05/10 01:36 A. Camsonne Timing shift possibly due to trigger change for efficiency run
  • 342824 12/05/10 01:33 Hisham TDC signals looks OK. run 9263
  • 342821 12/05/10 01:17 hkkang scaler log for 9262
  • 342820 12/05/10 01:16 hkkang end-of-run 9262
  • 342819 12/05/10 01:02 hisham TDC signals Shift
  • 342816 12/05/10 00:36 hkkang scaler info of 9261
  • 342815 12/05/10 00:35 hkkang end-of-run 9261
  • 342812 12/05/10 00:00 hyao Swing Shift Summary
  • Saturday

  • 342810 12/04/10 23:43 hyao target pressure and power graph
  • 342809 12/04/10 23:42 hyao Target Tools screen
  • 342807 12/04/10 23:13 eschulte End of Run 9259 CODA
  • 342806 12/04/10 23:12 eschulte End of Run 9259 Scalers
  • 342805 12/04/10 23:12 eschulte End of Run 9259 ROCS
  • 342804 12/04/10 23:11 eschulte End of Run 9258 scalers
  • 342803 12/04/10 22:21 Carlos Golden run updated to 9258
  • 342801 12/04/10 22:12 eschulte End of Run 9258 CODA
  • 342799 12/04/10 22:11 eschulte End of Run 9258 ROCS
  • 342798 12/04/10 21:41 eschulte reset methane alarm
  • 342797 12/04/10 21:31 eschulte Reset methane gas alarm
  • 342796 12/04/10 21:17 eschulte Should be run 9257
  • 342794 12/04/10 21:12 eschulte End of Run 9275 CODA
  • 342792 12/04/10 21:12 eschulte End of Run 9257 scalers
  • 342791 12/04/10 21:11 eschulte End of Run 9257 ROCS
  • 342790 12/04/10 20:57 eschulte Calorimeter Currents
  • 342789 12/04/10 20:53 hyao Beam permit and take data
  • 342788 12/04/10 20:19 A. Camsonne HRS timing check
  • 342787 12/04/10 20:16 eschulte Updated Database
  • 342785 12/04/10 19:54 eschulte Restarted CODA
  • 342776 12/04/10 18:41 hyao LHRS P=2.962 GeV for kin 2
  • 342775 12/04/10 18:27 hyao Escorted Access
  • 342773 12/04/10 18:18 eschulte End of Run 9253 CODA
  • 342772 12/04/10 18:18 eschulte End of Runm 9253 scalers
  • 342771 12/04/10 18:17 eschulte End of Run 9253 ROCS
  • 342769 12/04/10 17:18 eschulte End of Run 9252 CODA
  • 342767 12/04/10 17:17 eschulte End of Run 9252 scalers
  • 342766 12/04/10 17:16 eschulte End of Run 9252 ROCS
  • 342763 12/04/10 16:14 eschulte End of Run 9251 CODA
  • 342762 12/04/10 16:14 eschulte End of Run 9251 scalers
  • 342761 12/04/10 16:12 eschulte End of Run 9251 ROCS
  • 342759 12/04/10 16:01 bogdanw shift summary
  • 342758 12/04/10 15:21 igorko target info
  • 342755 12/04/10 15:10 Lindgren End of Run 9250
  • 342754 12/04/10 15:08 LINDGREN ENEND OF RUN 9250 SCALERS
  • 342753 12/04/10 14:11 igorko Calo anode currents
  • 342750 12/04/10 14:08 Lindgren End of Run 9249
  • 342749 12/04/10 14:07 A. Camsonne Updated ROC3 pedestal file
  • 342748 12/04/10 14:03 Lindgren SCALERS END OF RUN 9249
  • 342747 12/04/10 13:08 igorko target info
  • 342746 12/04/10 13:01 Lindgren IOC alarm
  • 342743 12/04/10 12:56 igorko end of run# 9248
  • 342740 12/04/10 12:00 igorko end of run# 9247
  • 342739 12/04/10 11:59 igorko end of run # 9247
  • 342738 12/04/10 11:57 igorko scalers info run# 9247
  • 342737 12/04/10 11:46 adaq s2 right ADC amplitude run #9247
  • 342729 12/04/10 10:49 igorko end of run #9246
  • 342728 12/04/10 10:48 igorko scalers info run# 9246
  • 342727 12/04/10 10:01 igorko Target alarm
  • 342726 12/04/10 09:47 igorko Target screenshots
  • 342722 12/04/10 09:08 igorko end of run #9244
  • 342721 12/04/10 09:06 igorko end of run #9244
  • 342719 12/04/10 09:03 igorko scalers info run #9244
  • 342718 12/04/10 08:50 adaq Gas Sensor alarm
  • 342717 12/04/10 08:19 R. Michaels re: writing to adaql1
  • 342713 12/04/10 08:00 efuchey 3-12-2010 Owl shift summary
  • 342712 12/04/10 07:59 hkkang scaler run9243
  • 342711 12/04/10 07:58 hkkang end-of-run 9243
  • 342710 12/04/10 07:32 M Friend JT valve stepped open
  • 342707 12/04/10 06:59 hkkang scaler run9242
  • 342706 12/04/10 06:58 hkkang end-of-run 9242
  • 342705 12/04/10 06:48 M Friend Target screenshots
  • 342704 12/04/10 06:34 efuchey Hall A tools
  • 342701 12/04/10 05:55 hkkang scaler info 9241
  • 342700 12/04/10 05:54 hkkang end-of-run 9241
  • 342697 12/04/10 05:23 hkkang end-of-run 9240
  • 342696 12/04/10 05:22 hkkang scaler for 9240
  • 342695 12/04/10 04:43 M Friend JT valve stepped open
  • 342692 12/04/10 04:24 hkkang scaler info run9239
  • 342691 12/04/10 04:23 hkkang end-of-run 9239
  • 342690 12/04/10 04:23 M Friend Calorimeter current snapshot
  • 342683 12/04/10 03:23 hkkang scaler of 9238
  • 342682 12/04/10 03:22 hkkang end-of-run 9238
  • 342681 12/04/10 03:00 M Friend Target low pressure alarm
  • 342680 12/04/10 02:36 Carlos Rates writing to adaql2
  • 342679 12/04/10 02:30 Carlos Back to /adaql2 only
  • 342678 12/04/10 02:28 Carlos Rates writing to adaql1
  • 342674 12/04/10 02:21 hkkang scaler info run9237
  • 342673 12/04/10 02:21 hkkang end-of-run 9237
  • 342672 12/04/10 02:20 M Friend Target screen snapshots
  • 342671 12/04/10 02:20 efuchey cryo flow
  • 342670 12/04/10 02:05 Carlos Rates writing to adaql1
  • 342669 12/04/10 01:50 Carlos Writing to adaql1
  • 342666 12/04/10 01:22 hkkang scaler log for 9236
  • 342665 12/04/10 01:21 hkkang end-of-run 9236
  • 342661 12/04/10 00:18 hkkang scaler info run9234
  • 342660 12/04/10 00:17 hkkang end-of-run 9234
  • Friday

  • 342658 12/03/10 23:59 hyao Swing Shift Summary
  • 342656 12/03/10 23:45 eschulte Run 9234 Efficiency Run
  • 342654 12/03/10 23:43 eschulte End of Run 9233 CODA
  • 342653 12/03/10 23:43 eschulte End of Run 9233 scalers
  • 342651 12/03/10 23:42 eschulte End of Run 9233 (really) ROCS
  • 342650 12/03/10 23:32 eschulte Jumped the gun...Not EOR 9233
  • 342649 12/03/10 23:29 eschulte End of Run 9233 ROCS
  • 342647 12/03/10 22:43 eschulte End of Run 9232 CODA
  • 342645 12/03/10 22:43 eschulte End of Run 9232 scalers
  • 342644 12/03/10 22:42 eschulte End of Run 9232 ROCS
  • 342641 12/03/10 21:43 eschulte End of Run 9231 CODA
  • 342640 12/03/10 21:42 eschulte End of Run 9231 Scalers
  • 342639 12/03/10 21:41 eschulte End of Run 9231 ROCS
  • 342638 12/03/10 21:06 A. Camsonne Summary test latest trigger firmware
  • 342636 12/03/10 20:39 A. Camsonne Trigger production firmware loaded
  • 342634 12/03/10 20:35 A. Camsonne Run 9230 with trigger - plots
  • 342632 12/03/10 20:30 A. Camsonne Run 9230 - ARS valid window set to 0x50
  • 342627 12/03/10 20:20 A. Camsonne Run 9228 with threshold to 0x650
  • 342625 12/03/10 20:16 xiaohui Target Screen
  • 342624 12/03/10 20:11 A. Camsonne Data with new trigger firware Run 9227
  • 342623 12/03/10 20:07 A. Camsonne Scalers rates Run 9227 with new trigger firmware
  • 342621 12/03/10 20:04 A. Camsonne Run with new Trigger firmware Run 9227
  • 342618 12/03/10 19:57 A. Camsonne New firmware loaded for trigger
  • 342617 12/03/10 19:55 eschulte End of Run 9225 CODA
  • 342615 12/03/10 19:54 eschulte End of Run 9225 Scalers
  • 342614 12/03/10 19:53 eschulte end of Run 925
  • 342608 12/03/10 19:14 A. Camsonne L1A accept with and without Compton running
  • 342606 12/03/10 18:55 eschulte End of Run 9224 CODA
  • 342604 12/03/10 18:54 eschulte End of Run 9224 ROCS
  • 342603 12/03/10 18:54 eschulte End of Run 9224 Scalers
  • 342601 12/03/10 17:56 eschulte End of Run 9223 CODA
  • 342599 12/03/10 17:54 eschulte End of Run 9223 Scalers
  • 342598 12/03/10 17:54 eschulte End of Run 9223 ROCS
  • 342597 12/03/10 17:22 eschulte plots from run 9223
  • 342596 12/03/10 17:22 eschulte CALO TDC's
  • 342595 12/03/10 17:20 eschulte more TDC plots
  • 342594 12/03/10 17:18 eschulte Left Arm Timing?
  • 342593 12/03/10 16:59 A. Camsonne Bad Compton tune
  • 342591 12/03/10 16:54 eschulte End of Run 9222 CODA
  • 342589 12/03/10 16:53 eschulte End of Run 9222 Scalers
  • 342588 12/03/10 16:51 eschulte Run 9222 End of Run
  • 342587 12/03/10 16:09 Carlos Straight beam
  • 342586 12/03/10 16:03 horn Day shift summary
  • 342584 12/03/10 15:59 Carlos Rates with straight tune
  • 342583 12/03/10 15:57 Carlos Rates with Compton tune
  • 342580 12/03/10 15:02 hisham Target snapshots
  • 342579 12/03/10 14:59 Carlos Trigger rates without Compton DAQ running
  • 342576 12/03/10 14:51 hisham Run 9220 Runcontrol, prescales, roc's
  • 342575 12/03/10 14:50 hisham Run 9220 scalers
  • 342574 12/03/10 14:40 Carlos Compton run stopped at 2:40pm
  • 342573 12/03/10 14:08 A. Camsonne Correlation between adaql2 dead time fluctuations and Compton DAQ
  • 342570 12/03/10 13:50 hisham run 9219 RunControl, prescales, rocs
  • 342569 12/03/10 13:48 hisham run 9219 scalers
  • 342566 12/03/10 12:43 hisham Target charts at the begining of hall C Qweak test
  • 342563 12/03/10 12:17 hisham Run92167 rocs
  • 342562 12/03/10 12:17 horn Calorimeter currents screenshot
  • 342561 12/03/10 12:16 hisham Run 6217 Runcontrol, prescales
  • 342560 12/03/10 12:14 hisham Run 9217 scalers
  • 342553 12/03/10 11:15 Hisham Run 9216 rocs
  • 342552 12/03/10 11:15 hisham Run 9216 RunControl, prescales
  • 342551 12/03/10 11:14 Hisham Run 9216 scalers
  • 342550 12/03/10 10:28 horn Beam positions adjusted
  • 342547 12/03/10 10:18 Carlos Rates change
  • 342546 12/03/10 10:14 Hisham Run 9215: rocs
  • 342545 12/03/10 10:13 Hisham Run 9215: Runcontrol, prescales
  • 342544 12/03/10 10:11 hisham run 9215 scalers
  • 342543 12/03/10 09:31 Hisham Fast feedback reset
  • 342540 12/03/10 09:19 Hisham run 9214 rocs
  • 342539 12/03/10 09:18 Hisham Run 9214 RunControl, prescales
  • 342538 12/03/10 09:16 hisham run 9214 scalers
  • 342537 12/03/10 08:19 horn Beam back
  • 342534 12/03/10 07:58 V. Sulkosky Owl Shift Summary, December 3
  • 342533 12/03/10 07:54 V. Sulkosky Beam Down
  • 342532 12/03/10 07:50 hoyoung Target screens
  • 342530 12/03/10 07:36 hkkang scaler log run9213
  • 342529 12/03/10 07:35 hkkang end-of-run 9213
  • 342526 12/03/10 06:52 hkkang scaler info 9212
  • 342525 12/03/10 06:51 hkkang end-of-run 9212
  • 342524 12/03/10 06:05 hoyoung Target screens
  • 342521 12/03/10 05:49 hkkang scaler log for 9211
  • 342520 12/03/10 05:48 hkkang end-of-run 9211
  • 342518 12/03/10 04:49 V. Sulkosky Re: Detector Flammable Gas Alarm
  • 342516 12/03/10 04:48 hkkang sclaer log for 9210
  • 342515 12/03/10 04:47 hkkang end-of-run 9210
  • 342514 12/03/10 04:31 V. Sulkosky Detector Flammable Gas Alarm
  • 342513 12/03/10 04:15 hoyoung Target screens
  • 342510 12/03/10 03:47 hkkang scaler log run9209
  • 342509 12/03/10 03:46 hkkang end-of-run 9209
  • 342503 12/03/10 02:57 V. Sulkosky Hall A Tools
  • 342500 12/03/10 02:44 hkkang scaler log for 9208
  • 342499 12/03/10 02:43 hkkang end-of-run 9208
  • 342498 12/03/10 02:20 V. Sulkosky Beam Stability
  • 342497 12/03/10 02:01 hoyoung Target screens
  • 342494 12/03/10 01:46 hkkang 9207 scaler info
  • 342493 12/03/10 01:45 hkkang scaler info 9207
  • 342492 12/03/10 01:44 hkkang end-of-run 9207
  • 342491 12/03/10 01:33 V. Sulkosky Disks on adaql2 Slowly Filling Up
  • 342490 12/03/10 01:28 V. Sulkosky Calorimeter DC Levels
  • 342487 12/03/10 00:45 hkkang scaler 9206
  • 342486 12/03/10 00:45 hkkang end-of-run 9206
  • 342485 12/03/10 00:36 hoyoung Low 15K flow
  • 342483 12/03/10 00:03 jpchen Swing shift summary
  • 342482 12/03/10 00:01 hoyoung Target screens
  • Thursday

  • 342479 12/02/10 23:30 jpchen bleedthough measurement
  • 342462 12/02/10 18:40 efuchey Compton background rates
  • 342461 12/02/10 18:25 efuchey Comparison between 1625 and 1650 V for compton detector
  • 342457 12/02/10 18:10 efuchey cavity beam scan
  • 342456 12/02/10 18:09 A. C Calorimeter run list with new firmware
  • 342453 12/02/10 17:53 efuchey Successful compton tune
  • 342452 12/02/10 17:49 A. Camsonne Removed discriminator after modification to logic module
  • 342449 12/02/10 16:47 efuchey Compton rates for 12-02
  • 342441 12/02/10 16:20 A. Camsonne Production firmware loaded in trigger
  • 342436 12/02/10 16:03 A. Deur day shift summary
  • 342427 12/02/10 15:31 A. Camsonne DAQ work for DVCS
  • 342414 12/02/10 14:38 A. Deur target snapshot +bad sensor in loop 2
  • 342396 12/02/10 11:57 A. Camsonne Calorimeter LED scan - run 9173
  • 342395 12/02/10 11:56 A. Camsonne ARS stop going through Lecroy discriminator
  • 342381 12/02/10 09:36 A. Camsonne Cosmics paddles on - HV increased to 800 V for trigger cosmics tests
  • 342377 12/02/10 09:15 efuchey Compton beamline modified
  • 342376 12/02/10 09:10 A. Deur target snapshot
  • 342373 12/02/10 08:56 A. Camsonne Trigger FPGA firmware loaded
  • 342368 12/02/10 08:11 hkkang scaler info 9163
  • 342367 12/02/10 08:10 hkkang end-of-run 9163
  • 342365 12/02/10 08:05 n.muangma Owl shift summary
  • 342364 12/02/10 07:52 n.muangma beam chart
  • 342363 12/02/10 07:51 n.muangma calorimeter
  • 342362 12/02/10 07:48 hoyoung Target screens
  • 342359 12/02/10 07:40 hkkang scaler info 9162
  • 342358 12/02/10 07:39 hkkang end-of-run 9162
  • 342357 12/02/10 07:31 fansler Nitrogen liquid levels
  • 342354 12/02/10 06:41 hkkang scaler log run9161
  • 342353 12/02/10 06:40 hkkang end-of-run 9161
  • 342352 12/02/10 06:31 Carlos Back to /adaql2 only
  • 342351 12/02/10 06:04 hoyoung Target screens
  • 342348 12/02/10 05:38 hkkang scaler log of 9160
  • 342347 12/02/10 05:37 hkkang end-of-run 9160
  • 342344 12/02/10 04:40 hkkang scaler 9159
  • 342343 12/02/10 04:39 hkkang end-of-run 9159
  • 342342 12/02/10 03:57 hoyoung Target screens
  • 342339 12/02/10 03:33 hkkang scaler log for 9158
  • 342338 12/02/10 03:33 hkkang end-of-run 9158
  • 342330 12/02/10 02:34 hkkang scaler info 9157
  • 342329 12/02/10 02:33 hkkang end-of-run 9157
  • 342328 12/02/10 02:04 hoyoung Target screens
  • 342327 12/02/10 01:37 Carlos Writing to adaql1 again
  • 342325 12/02/10 01:28 n.muangma long FSD trip
  • 342324 12/02/10 01:15 Carlos Run 9156: high DVCS-clk rate due to IOC reboot
  • 342323 12/02/10 01:15 n.muangma ioc reboot
  • 342321 12/02/10 01:13 hkkang scaler info 9156
  • 342320 12/02/10 01:11 hkkang screenshot-9156
  • 342317 12/02/10 00:31 hkkang scaler info 9155
  • 342316 12/02/10 00:30 hkkang end-of-run9155
  • 342315 12/02/10 00:03 hoyoung Target screens
  • Wednesday

  • 342313 12/01/10 23:58 Xiaohui Zhan Swing Shift Summary
  • 342310 12/01/10 23:26 A. Tobias ROCs for Efficiency Run 9154
  • 342309 12/01/10 23:25 A. Tobias RunControl and prescales for Efficiency Run 9154
  • 342308 12/01/10 23:24 A. Tobias Scalars for run 9154
  • 342307 12/01/10 23:23 Xiaohui Zhan DVCS ADC screen
  • 342306 12/01/10 23:21 hyao target pressure and power screen
  • 342303 12/01/10 22:51 A. Tobias ROCs for run 9153
  • 342302 12/01/10 22:50 A. Tobias RunControl and prescales for run 9153
  • 342301 12/01/10 22:49 A. Tobias Scalars for run 9153
  • 342300 12/01/10 22:20 Xiaohui Zhan BPMA Y position drifted to 0.4 between 21:45 to 22:12
  • 342299 12/01/10 22:15 A. Tobias BPMA y position moved
  • 342298 12/01/10 22:04 Carlos adaql1 writing switched off
  • 342295 12/01/10 21:53 A. Tobias ROCs for run 9152: Deadtime seems to be higher compared to earlier in the shift
  • 342294 12/01/10 21:51 A. Tobias RunControl and prescales for run 9152
  • 342293 12/01/10 21:50 A. Tobias Scalars for run 9152
  • 342292 12/01/10 21:29 hyao target screen
  • 342289 12/01/10 20:52 A. Tobias ROCs for run 9151
  • 342288 12/01/10 20:51 A. Tobias RunControl and prescales for run 9151
  • 342287 12/01/10 20:50 A. Tobias Scalars for run 9151
  • 342285 12/01/10 20:01 Carlos DVCS IOC going bad
  • 342284 12/01/10 19:56 Xiaohui Zhan DVCS IOC rebooted
  • 342282 12/01/10 19:47 A. Tobias ROCs for run 9150
  • 342281 12/01/10 19:47 A. Tobias RunControl and prescales for run 9150
  • 342280 12/01/10 19:46 A. Tobias Scalars for run 9150
  • 342272 12/01/10 18:45 A. Tobias ROCs for run 9149
  • 342271 12/01/10 18:44 A. Tobias RunControl and prescales for run 9149
  • 342270 12/01/10 18:43 A. Tobias Scalars for run 9149
  • 342269 12/01/10 18:26 A. Camsonne Turned adaql1 writing on
  • 342266 12/01/10 17:51 A. Tobias ROCs for run 914
  • 342265 12/01/10 17:50 A. Tobias RunControl & prescales for run 914
  • 342264 12/01/10 17:49 A. Tobias Scalars for run 9148
  • 342263 12/01/10 17:12 Carlos Alarm set on DVCS IOC (iochacs)
  • 342261 12/01/10 16:39 A. Tobias Ended run 9147 early due to no beam
  • 342260 12/01/10 16:37 A. Tobias ROCs for run 9147 (oops, I took it after run ended)
  • 342259 12/01/10 16:36 Xiaohui Zhan MCC stopped beam for dispersion check
  • 342258 12/01/10 16:35 A. Tobias RunControl and prescales for run 9147
  • 342255 12/01/10 15:56 hyde Shift Summary Day 01 December 2010
  • 342252 12/01/10 15:52 hyde run 9146 scalers
  • 342251 12/01/10 15:51 hyde roc run 9146
  • 342248 12/01/10 15:05 zana Rocs and RunControl Run 9145
  • 342247 12/01/10 15:03 zana Scaler Run 9145
  • 342246 12/01/10 14:36 igorko target status
  • 342245 12/01/10 14:34 hyde DVCS reboot proceedure
  • 342244 12/01/10 14:29 hyde led Clock and DVCS IOC reboot
  • 342243 12/01/10 14:23 hyde DVCS ADC
  • 342242 12/01/10 14:19 hyde Run 9145 clk rate fluctuation
  • 342240 12/01/10 14:15 hyde reboot DVCS IOC
  • 342239 12/01/10 14:07 zana ocs and RunControl run 9144
  • 342238 12/01/10 14:06 zana Scaler Run 9144
  • 342237 12/01/10 13:56 hyde Beam without Hall C
  • 342236 12/01/10 13:32 hyde Hall C off 20 min
  • 342234 12/01/10 13:06 hyde N2 levels HRS-L
  • 342232 12/01/10 13:06 zana Rocs and RunControl run 9143
  • 342231 12/01/10 13:04 zana Scaler run 9143
  • 342230 12/01/10 12:38 hyde DVCS Calo DC Run 9143
  • 342229 12/01/10 12:32 hyde HRS-L N2 Cryo level
  • 342228 12/01/10 12:20 hyde Run 9127 deadtime all splits
  • 342227 12/01/10 12:17 A. Camsonne Trigger readout empty
  • 342226 12/01/10 12:11 hyde Afterpusle on DVtoF x10
  • 342225 12/01/10 12:08 igorko target status
  • 342221 12/01/10 12:03 hyde BCM VtoFx10 signals run 9140
  • 342211 12/01/10 11:28 hyde left/right bcm
  • 342210 12/01/10 11:03 hyde Deadtime Run 9139
  • 342207 12/01/10 10:54 zana Rocs and RunControl run 9139
  • 342206 12/01/10 10:53 zana Scaler run 8139
  • 342205 12/01/10 10:18 hyde Run 9138 deadtime
  • 342204 12/01/10 10:02 Igor Korover target screen
  • 342201 12/01/10 09:52 zana Rocs and RunControl run 9138
  • 342200 12/01/10 09:51 zana Scaler run 9138
  • 342199 12/01/10 09:34 hyde Run 9138 unstable beam
  • 342198 12/01/10 09:08 hyde Run 9121 optics Cer Efficiency
  • 342197 12/01/10 08:54 hyde Half Wave Plate moved to OUT
  • 342194 12/01/10 08:50 zana Rocs and RunControl run 9137
  • 342193 12/01/10 08:48 zana scaler run 9137
  • 342190 12/01/10 08:21 zana rocs and runControl run 9136
  • 342189 12/01/10 08:20 zana Scaler run 9136
  • 342188 12/01/10 08:02 n.muangma Owl shift summary
  • 342186 12/01/10 07:37 hoyoung Target screens
  • 342183 12/01/10 07:18 hkkang scaler of 9135
  • 342182 12/01/10 07:17 hkkang end-of-run 9135
  • 342179 12/01/10 06:18 hkkang scaler of 9134
  • 342178 12/01/10 06:17 hkkang end-of-run 9134
  • 342177 12/01/10 05:50 hoyoung Target screens
  • 342176 12/01/10 05:48 n.muangma beam trip
  • 342173 12/01/10 05:14 hkkang scaler 9133
  • 342172 12/01/10 05:13 hkkang end-of-run summary 9133
  • 342169 12/01/10 04:36 hkkang scaler 9132
  • 342168 12/01/10 04:35 hkkang run summary 9132
  • 342167 12/01/10 03:37 hoyoung Target screens
  • 342165 12/01/10 03:33 hkkang scaler of 9131
  • 342164 12/01/10 03:33 hkkang end-of-run 9131
  • 342156 12/01/10 02:33 hkkang scaler log 9130
  • 342155 12/01/10 02:32 hkkang end-of-run 9130
  • 342154 12/01/10 01:41 hoyoung Target screens
  • 342151 12/01/10 01:31 hkkang scaler 9129
  • 342150 12/01/10 01:31 hkkang end-of-run 9129
  • 342147 12/01/10 00:28 hkkang scaler panel for 9128
  • 342146 12/01/10 00:26 hkkang end-of-run screen shot